### v2

# Orthogonal projections of hypercubes

Abstract
Projections of hypercubes are widely used for visualizing high-dimensional binary state spaces across various scientific fields. Conventional methods for projecting hypercubes, however, face practical difficulties. Manual methods require nontrivial adjustments of the projection basis, while optimization-based algorithms limit the interpretability and reproducibility of the resulting plots. These limitations motivate the exploration of theoretically analyzable projection algorithms such as principal component analysis (PCA). Here, we investigate the mathematical properties of PCA for projecting hypercubes. Our numerical and analytical results show that PCA effectively captures polarized distributions within the state space. This property enables the assessment of the asymptotic distribution of projected vertices and error bounds, thus characterizing the performance of PCA in the projected space. We demonstrate the application of PCA to visualize the hypercubic energy landscapes of Ising spin systems. By adding projected hypercubic edges, these visualizations reveal pathways of correlated spin flips. Our work contributes to a better understanding of how PCA can be applied to explore high-dimensional binary data.
I Introduction
Human interest in higher dimensions, particularly the fourth dimension, has been emerging since the 19th century [1, 2, 3, 4]—before the theory of relativity by A. Einstein [5, 6], which is well known as the theory of four-dimensional spacetime [7]. Unlike the theory of relativity, 19th-century mathematicians, physicists, and philosophers focused on the spatial fourth dimension. Among them, C. H. Hinton has been influential because his works [8] indicate that the tesseract, or four-dimensional cube, can embody four-dimensional space. Since then high-dimensional space has been explored with hypercubes.
By generalizing the idea of a three-dimensional cube to higher dimensions, one obtains a high-dimensional cube or a hypercube [9] (see Appendix A for details on the construction of a hypercube). Although a hypercube is purely a geometric concept [10, 11, 9, 12], it has applications not only in physics [13, 14, 15], electrical engineering [16, 17, 18, 19, 20, 21], graph theory [22, 17], and recreational mathematics [17, 12, 23], but also in a wide range of interdisciplinary fields [24]. In various arts, e.g., literature [25, 26, 23], visual art [27, 1, 28, 29], architecture [27], ornament [30], and film [31, 32, 33, 34], hypercubes have repeatedly inspired human imagination and have become an embodiment of high-dimensional space. Since the pioneering work of almost 150 years ago [10] and subsequent contributions [8, 9], interest in high-dimensional geometry, especially hypercubes, has been continuously growing.
Returning to the sciences, hypercubes visualize binary state111 Also called microstate, configuration, or phase. space in diverse fields, ranging from physics to biology.222 See, for instance, evolutionary landscapes [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], epistasis [35, 47, 48], chemical reaction networks [49, 50], learning in neural networks [51], genetic code space [52, 53, 44], allostery [54, 55, 56, 57, 58, 59], quantum states [60, 61], quantum walks [62], language space [63], data visualization [64, 65], probability currents [66], protein folding [67, 68], energy landscapes [69], gene regulation [70, 71, 72, 73, 74], quantum many-body scar states [75, 76, 77], morphology [78, 79], disease progression [44, 45], gene regulatory networks [80], gene expression [81], and spike dynamics [82, 83]. This is because hypercubic vertices correspond to binary states of a system [Figs. 1(a) and 1(b)], and hypercubic edges represent transitions333 Strictly speaking, only one of the dimensions, components, or elements of the system must change in the transition on the hypercubic edge. This kind of dynamics is called asynchronous dynamics (update) [84, 85, 86, 80]. It is also called multipartite dynamics [87, 88, 82, 83] in the context of stochastic thermodynamics. On the other hand, transitions with several changes of elements are called synchronous dynamics (update) [89, 90, 91, 92, 93, 94, 95]. between states [Figs. 1(c) and 1(d)]. This can be naturally applied to illustrate the high-dimensional binary state space or state transition diagram. We refer to such a binary state space as a hypercubic state space. As an example of transitions between the vertices, we visualize the Hamiltonian path on a hypercube [Figs. 1(c) and 1(d), Tables 1 and 2].
By using these useful two-dimensional illustrations of hypercubes, scientists have visualized high-dimensional hypercubic state spaces to intuitively understand the state space structure and state transition dynamics of systems of interest. Despite the practical usefulness of hypercubes, in general, it is “disappointingly difficult” to visualize hypercubes, even for low-dimensional ones [96]. Thus, studies on projecting the hypercube have targeted relatively lower-dimensional systems, and the procedure to project hypercubic state spaces is still unclear.
One method to project a hypercubic state space, or a hypercube, is orthogonal projection [9], which is reproducible. In orthogonal projection, by casting a shadow perpendicular to the two-dimensional plane with a distant light source, the parallels and lengths are preserved between the edges representing the same dimensions. Despite this useful property, the weakness of orthogonal projection is that one must manually determine the alignment of the object to the targeted plane: the projection of the unit vector of each dimension is manually determined. For projections of higher-dimensional hypercubes, it is nontrivial and impractical to manually determine the projection of the unit vectors. How can one decide the projection of the unit vector of each dimension to create informative projections of hypercubes?
Figure 1: A gallery of hypercubes and Hamiltonian (directed) paths on them. (a) A cube, or three-dimensional hypercube, with three-dimensional coordinates of vertices. (b) A tesseract, or four-dimensional hypercube, with four-dimensional coordinates of vertices. (c) A Hamiltonian (directed) path on a cube. (d) A Hamiltonian (directed) path on a tesseract. In (c) and (d), the arrow indicates the direction of the path. One obtains a Hamiltonian path by converting the decimals to Gray code [16, 12] and following them in ascending order. See Tables 1 and 2 for Gray codes used to visualize Hamiltonian paths on three-dimensional and four-dimensional hypercubes.
Decimal Binary Gray
000 000 000
001 001 001
002 010 011
003 011 010
004 100 110
005 101 111
006 110 101
007 111 100
Table 1: Decimal identifiers of three-dimensional hypercubic vertices, and corresponding binary and Gray codes [16, 17]. The first digit (left-most digit) of the Gray code is the same as the binary code. One obtains the digit of the Gray code by performing an exclusive-or (xor) operation on the corresponding digit of the binary code with its left digit.
Decimal Binary Gray
$0$ $0$ $0$
0001 0001 0001
0002 0010 0011
0003 0011 0010
0004 0100 0110
0005 0101 0111
0006 0110 0101
0007 0111 0100
0008 1000 1100
0009 1001 1101
0010 1010 $1$
0011 1011 1110
0012 1100 1010
0013 1101 1011
0014 1110 1001
0015 $1$ 1000
Table 2: Decimal identifiers of four-dimensional hypercubic vertices, and corresponding binary and Gray codes [16, 17]. The first digit (left-most digit) of the Gray code is the same as the binary code. One obtains the digit of the Gray code by performing an exclusive-or (xor) operation on the corresponding digit of the binary code with its left digit.
An alternative method to project a hypercube is optimizing the coordinates of projected vertices following a predefined error function [96]. While this method projects hypercubes automatically, it has several limitations. The first limitation is reproducibility. Because of the stochastic nature of the optimization process, projections generated from the same vertices are not guaranteed to be the same (unless all relevant parameters are fixed). Interpretability is another weakness of projecting hypercubes by optimization. Because of its nonlinear nature, what is indicated from the resulting plot is not obvious (see, however, [97, 98, 99]). How can one create reproducible and interpretable projections of hypercubes?
Here, we reveal that principal component analysis (PCA) [100, 101] provides reproducible, interpretable, and automatic projections of hypercubes. PCA is a linear dimensionality reduction method frequently performed in statistics and machine learning [102, 103]. By interpreting the principal component (PC) loading as the basis for the projection of a hypercube, we show that one can draw the edges of a hypercube on a projected two-dimensional plane. This idea is closely related to the biplot [104], a method to visualize loadings (eigenvectors) with data points to assist in the interpretation of resulting plots. By combining ideas from geometry and statistics (machine learning), we achieve informative projections of hypercubes [78]: we can analytically obtain the properties of the projection, and the resulting plots are examined with such prior knowledge.
This paper is organized as follows. In Sec. II, we present the Hamming and fractal projections with introducing manual orthogonal projections of hypercubes. Then, in Sec. III, we show how biplots enable us to interpret the resulting plots of PCA on hypercubic vertices. We provide some examples of orthogonal projections of hypercubes using PCA in Sec. IV. Through analytical and numerical investigations in Sec. V, we show that principal components are informative for weighted vertices. In Sec. VI, using the inner-product error, we show that the vertices around the origin of the projected space less accurately preserve the original distances between them. We apply our method to visualize the hypercubic energy landscape [69] of Ising spin systems in Sec. VII, and conclude this paper in Sec. VIII.
II Orthogonal Projections of Hypercubes
Orthogonal projection is a linear method to project a high-dimensional object. We introduce the concept of the contribution basis, which is the projected unit vector of each dimension. We then illustrate isometric, Hamming, and fractal projections of hypercubes by manually varying the contribution basis.
Figure 2: A gallery of orthogonal projections of hypercubes. Colored arrows represent the contribution basis corresponding to the original dimensions. The boxes on the bottom right indicate the correspondence between the colors and the original dimensions. (a) An isometric projection of a cube. Notice that $\begin{bmatrix}1&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&0\end{bmatrix}^{\top}$ are overlapped. (b) An isometric projection of a tesseract. (c) A Hamming projection of a cube. (d) A Hamming projection of a tesseract. Notice that $\begin{bmatrix}1&0&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&1&0\end{bmatrix}^{\top}$ are overlapped. In Hamming projections (c) and (d), the contribution basis of each dimension has the same horizontal contribution. (e) A fractal projection of a six-dimensional hypercube. (f) A fractal projection of an eight-dimensional hypercube. In fractal projections (e) and (f), the contribution basis of the first half (left half) of the code is ten times longer than the rest. More projections of hypercubes are available in the Supplemental Material [105].
II.1 Dimensionality reduction and orthogonal projection
To visualize hypercubic vertices or high-dimensional binary data, one needs to reduce the dimensionality of the data. Dimensionality reduction methods are roughly divided into two categories: nonlinear and linear methods. The former often employ optimization of target functions to determine the projected coordinates, while the latter corresponds to linear projection. Modern nonlinear dimensionality reduction methods, such as t-distributed stochastic neighbor embedding (t-SNE) [106] and uniform manifold approximation and projection (UMAP) [107], project high-dimensional objects by optimizing target functions with an emphasis on patterns or clusters in the data. Although the resulting plots preserve the neighboring relationships between vertices well, it is nontrivial to interpret the meaning of the projected coordinates or to extract information from the resulting plots. On the other hand, resulting plots of linear methods are more intuitive because of their linearity: the projection is a shadow of the object.
A linear method to project a high-dimensional object or polytope is suggested in high-dimensional geometry [9]. Parallel projection illustrates polytopes (or high-dimensional polyhedra) in lower-dimensional space by moving the vertices of the object parallel to the selected direction until they reach the desired lower-dimensional space. Orthogonal projection is a special form of parallel projection where the selected direction is orthogonal to the lower-dimensional space. It is known that orthogonal projection can be derived from concentrically overlapping the cross-sections (sections) of a polytope and connecting the pairs of vertices of edges, with the condition that the cross-sections include the vertices and are parallel to the targeted lower-dimensional space of the object. For projections of hypercubes, orthogonal projection is a natural choice: edges that are parallel before the projection remain parallel, and edges that are parallel to each other share the same length after the projection [108], which assists in the interpretation of the resulting plots.
The coordinates $\bm{r}\coloneqq\begin{bmatrix}r_{1}&r_{2}\end{bmatrix}^{\top}\in\mathbb{R}^{2}$ of an orthogonally projected hypercubic vertex in two-dimensional space are given as a linear transformation of binary coordinates $\bm{b}$:
[$\ \bm{r}\\ \vrule\end{bmatrix}=\begin{bmatrix}\text{---}&\bm{v}_{1}^{\top}&\text{---}\\ \text{---}&\bm{v}_{2}^{\top}&\text{---}\end{bmatrix}\begin{bmatrix}\vrule\\ \bm{b}\\ \vrule\end{bmatrix},$ (1)
where $\bm{b}\coloneqq\begin{bmatrix}b_{1}&\cdots&b_{N}\end{bmatrix}^{\top}\in\left\{1,0\right\}^{N}$ is an $N$-dimensional coordinate of the hypercubic vertices. The reduced original dimensions are represented by two $N$-dimensional vectors, $\bm{v}_{1},\bm{v}_{2}\in\mathbb{R}^{N}$.
II.2 Contribution basis
To understand the resulting projection, we rewrite [9] Eq. (1) using the contribution basis,
$\bm{r}=\sum_{i=1}^{N}b_{i}\bm{e}_{i},$ (2)
where the contribution basis,
$\bm{e}_{i}\coloneqq\begin{bmatrix}v_{i;1}\\ v_{i;2}\end{bmatrix},$ (3)
is the $i$th column of the projection matrix $\begin{bmatrix}\bm{v}_{1}&\bm{v}_{2}\end{bmatrix}^{\top}$,
[$\ \bm{e}_{1}&\cdots&\bm{e}_{N}\\ \vrule&&\vrule\end{bmatrix}\coloneqq\begin{bmatrix}\text{---}&\bm{v}_{1}^{\top}&\text{---}\\ \text{---}&\bm{v}_{2}^{\top}&\text{---}\end{bmatrix}.$ (4)
Here, $v_{i;j}$ is the $i$th element of $\bm{v}_{j}$.
The interpretation of the contribution basis in Eq. (2) is the contribution from the $i$th dimension to the projected space. Because $b_{i}\in\left\{1,0\right\}$, the contribution basis corresponds to the projected unit vector of each dimension. Thus, by introducing the contribution basis, we can decompose the resulting projection into the contributions from each dimension as shown in Eq. (2). The remaining question is how to determine the contribution bases $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$.
II.3 Isometric projection
One particularly regular and symmetric projection is called isometric projection, where a Petrie polygon is projected as a regular polygon located at the periphery of the projection of a polytope [9]. Petrie polygons are equatorial polygons, lying in planes crossing the center of the object and inscribed in great circles of the circumsphere of the object. To achieve isometric projection, one needs the projected unit vector of the $i$th dimension to form consecutive edges of the projected Petrie polygon, i.e., $\bm{e}_{i}$ form a Petrie polygon together with a reversed vector $-\bm{e}_{i}$. To project the Petrie polygon as a regular polygon, the contribution bases are determined as $\bm{e}_{i}=\begin{bmatrix}\cos\left(\theta_{i}\right)&\sin\left(\theta_{i}\right)\end{bmatrix}^{\top}$, where $\theta_{i}=\frac{\uppi}{N}\left(i-1\right)+\phi$, with a constant $\phi$. The resulting isometric projections and vectors $\bm{e}_{i}$ used to create them are shown in Figs. 2(a) and 2(b), where Petrie polygons are projected as a regular hexagon [Fig. 2(a)] and octagon [Fig. 2(b)]. We determine the angle $\phi=\frac{\uppi}{2N}$ for symmetry of the projections. In isometric projections, all edges of the hypercube are drawn with the same length.
II.4 Hamming and fractal projections
By modifying the contribution bases $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$ for projection, one can view a high-dimensional hypercube from various angles. Several methods reflecting the cross-section of the hypercube have already been suggested [9]. Here, we introduce two other orthogonal projections of hypercubes used in sciences.
A Hamming projection [Figs. 2(c) and 2(d)], we call, is a type of projection that plots vertices according to the Hamming distance [109] from a selected vertex. To achieve this, an example of a contribution basis is $\bm{e}_{i}=\begin{bmatrix}1&\left(i-\frac{N+1}{2}\right)\frac{2}{N-1}\sin\left(\frac{\varphi}{2}\right)\end{bmatrix}^{\top}$, where $\varphi\coloneqq\arccos\left(\frac{\bm{e}_{1}^{\top}\bm{e}_{N}}{\sqrt{\bm{e}_{1}^{\top}\bm{e}_{1}}\sqrt{\bm{e}_{N}^{\top}\bm{e}_{N}}}\right)$ is the angle between the first and last contribution basis, but it can be any. The first element of $\bm{e}_{i}$ is the same for all $i$. The second element can be any, but we determine them to increase linearly with $i$. This type of projection is found not only in research in physics444 Examples are found in quantum walk [62], unsupervised learning of states of Ising spin system [110, 111, 112], and quantum many-body scar states [75, 76, 77]. but also in biology.555 Examples include evolutionary landscapes [35, 39, 40, 41, 42, 43, 45, 46], epistasis [35, 47], genetic code space [44], data visualization [64, 65], protein folding [67, 68], gene regulation [70], disease progression [44, 45], and allostery [58]. The horizontal distance between the Hamming projected hypercubic vertices corresponds to the Hamming distance from a reference vertex on the rightmost or leftmost side of the projected space. In Figs. 2(c) and 2(d), the horizontal axis corresponds to the Hamming distance from the origin $\begin{bmatrix}0&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&0&0&0\end{bmatrix}^{\top}$, respectively. Notice that this projection corresponds to the Hasse diagram [113].
[…]
II.5 Exchanging the labels of vertices
[…]
Here, we show a visual understanding of this conversion through contribution bases. By reversing the direction of the contribution basis, one can swap the projected coordinates of the rest, cf. Figs. 2(a) and 2(b), or Figs. 3(a) and 3(b). The dimension of the reversed contribution basis corresponds to the dimension where there is a $1$ in $\bm{b}^{\prime}\oplus\bm{b}^{\prime\prime}$. Visualizing the projected vector provides a way to understand exchanging the labels via reversing the direction of unit vectors.
Figure 3: Exchanging the labels of vertices by reversing the direction of the contribution basis of corresponding digits. The boxes on the bottom right indicate the correspondence between the colors and the dimensions. (a) By reversing the contribution basis for the second digit, one can exchange the cubic labels of pairs of vertices, $\begin{bmatrix}0&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&0\end{bmatrix}^{\top}$, $\begin{bmatrix}1&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}1&1&0\end{bmatrix}^{\top}$, $\begin{bmatrix}0&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&1\end{bmatrix}^{\top}$, and $\begin{bmatrix}1&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}1&1&1\end{bmatrix}^{\top}$. Compare this with Fig. 2(a). (b) Obtaining different labels for a tesseract by swapping two contribution bases. The first and fourth unit vectors are reversed, cf. Fig. 2(b).
II.6 The limitation of manual orthogonal projections
Although one can create reproducible and interpretable (and even visually label-exchangeable) two-dimensional projections of hypercubes in the abovementioned way, the disadvantage is that one must manually determine the contribution basis of each dimension. For low-dimensional hypercubes, it is manageable to adjust the contribution basis oneself to make the projections easy to interpret. Still, it is nontrivial to decide on contribution bases for high-dimensional hypercubes. Thus, automatic and unsupervised projection methods are demanded, especially for projections of high-dimensional hypercubes. How can one project a hypercube with such a strategy?
III Orthogonal projections of hypercubes using PCA
To answer the question raised in the previous section—how can one create a linear but automatic and unsupervised projection of hypercubes—we suggest using a linear dimensionality reduction method: PCA. In this section, we derive PCA from minimizing an error function, but how can one interpret the projection of a hypercube using PCA? We show that biplot enable us to interpret PCA as an orthogonal projection of a hypercube.
III.1 PCA by minimizing inner-product error
Unsupervised and automatic dimensionality reduction is possible by minimizing an error function to create projections of high-dimensional objects. Unlike the previous study [96], where the error function is based on the difference between the pairwise Euclidean distance in low-dimensional projected space and the Hamming distance in high-dimensional original space, we suggest minimizing the difference of the inner product between vertices in the projected and original space. The inner product measures the similarity between two vectors, thus minimizing the difference between them results in a projection that preserves the original similarity between the vertices.
We define the error as the difference of the inner product in the original space and the projected space:
$\varepsilon\left(\bm{s},\bm{s}^{\prime}\right)\coloneqq\bm{s}^{\top}\bm{s}^{\prime}-\bm{r}^{\top}\left(\bm{s}\right)\bm{r}\left(\bm{s}^{\prime}\right),$ (5)
where, $\bm{s}\coloneqq\begin{bmatrix}s_{1}&\cdots&s_{N}\end{bmatrix}^{\top}\in\left\{+1,-1\right\}^{N}$ is a coordinate of the hypercubic vertices in the original space but with the Ising spin variable [114, 115, 116, 117]. Instead of binary variables $b_{i}\in\left\{1,0\right\}$, we introduce Ising spin variables $s_{i}=2b_{i}-1\in\left\{+1,-1\right\}$ to calculate the inner product because the inner product among the Ising variables reflects the similarity or overlap between the vertices. We need to find the vectors in the projection matrix $\begin{bmatrix}\bm{v}_{1}&\bm{v}_{2}\end{bmatrix}^{\top}$ of Eq. (1) that minimizes the mean squared error
$\varepsilon^{2}\right>\coloneqq\sum_{\bm{s},\bm{s}^{\prime}}p\left(\bm{s}\right)p\left(\bm{s}^{\prime}\right)\varepsilon^{2}\left(\bm{s},\bm{s}^{\prime}\right),$ (6)
where $p\left(\bm{s}\right)\in\left[0,1\right]$ is the normalized weight, i.e., the probability of finding the Ising coordinate $\bm{s}$, satisfying $\sum_{\bm{s}}p\left(\bm{s}\right)=1$.
Minimizing $\varepsilon^{2}\right>$ with the normalization constraint $\left\{\bm{v}_{i}^{\top}\bm{v}_{i}=1\right\}_{i=1}^{2}$ corresponds to classical multidimensional scaling (MDS) [118, 119]. Classical MDS666 Also called classical scaling, Torgerson scaling, Torgerson–Gower scaling, or principal coordinates analysis. with squared Euclidean (Pythagorean) distance provides lower-dimensional scaled coordinates preserving the original distance between the vertices. It is shown that classical MDS with squared Euclidean distance is equivalent to PCA [120], which is a linear dimensionality reduction method. See relevant studies [121, 122, 123] for the equivalence of classical MDS and PCA. Interested readers can obtain the derivation of PCA from the minimum inner-product error formulation and other formulations (maximum projection variance formulation and minimum reconstruction error formulation) in Appendix B.
The minimum inner-product error formulation of PCA has two advantages. First, unlike other formulations of PCA, the minimum inner-product error formulation minimizes a function that sums a term involving pairs of vertices [Eq. (6)] rather than minimizing (or maximizing) a function summing a term involving a single vertex [Eqs. (47) and (53)]. Thus, we can examine which distortions between pairs of vertices contribute to the error. We numerically and analytically investigate the quality of the projection of hypercubes through the inner-product error in Sec. VI by taking advantages of this feature. Second, this formulation is useful for understanding the projection of hypercubes because we can investigate where the error arises in the projection. The minimum inner-product error formulation of PCA can provide a visual understanding of distortion in the projection of hypercubes. Unlike the variance, reconstruction error or distance, the inner-product error can be readily estimated from the resulting plot. Through the minimum inner-product error formulation of PCA, we can visually estimate the quality of the projection of hypercubes.
As we show in the following Secs. V.1, VI.2 and VI.3, the projection of hypercubes using PCA is informative because the properties of the projection can be analytically obtained. In addition to the biplot-assisted interpretation of PCA in the next Sec. III.2, we can analytically obtain the properties of the projection of hypercubes for special cases. We analytically reveal both the strengths and weaknesses of PCA in projecting hypercubes in Secs. V.1, VI.2 and VI.3. With such prior knowledge of the projection of hypercubes, we can examine the resulting plots more deeply, even if the knowledge is from special cases.
III.2 PCA and biplot
Considering that PCA is an unsupervised and linear method for projection, we suggest interpreting PCA on binary vertices as a reproducible, interpretable, and automatic projection of a hypercube. PCA [100, 101], a method of statistics and unsupervised machine learning, is an essential technique for analyzing high-dimensional data [102, 103, 124, 125]. By calculating eigenvalues (explained variance) and eigenvectors (PC loadings) of the covariance matrix, PCA finds a set of vectors that maximize the variance of the data777 See Appendix B.1 for the derivation of PCA by the maximum projection variance formulation. and is often employed to perform linear dimensionality reduction.888 In many practical applications of PCA, the probability distribution $p\left(\bm{s}\right)$ is approximated by the empirical distribution of the data points. To perform PCA, one needs to calculate the covariance matrix of the data points. Specifically, in the context of hypercubic vertices, the covariance matrix $\bm{\varSigma}\in\mathbb{R}^{N\times N}$ is defined as
[…]
where $\bm{s}\right>\coloneqq\sum_{\bm{s}}p\left(\bm{s}\right)\bm{s}$ is the mean vector of the Ising coordinate.
Obtaining the eigenvalues $\left\{\lambda_{i}\right\}_{i=1}^{N}$ and eigenvectors $\left\{\bm{u}_{i}\right\}_{i=1}^{N}$ of the covariance matrix corresponds to finding the PCs of the hypercubic vertices. The eigenvalue equation is
$\bm{\varSigma}\bm{u}_{i}=\lambda_{i}\bm{u}_{i},$ (8)
[…]
$r_{i}\left(\bm{s}\right)=\bm{u}_{i}^{\top}\bm{s}.$ (9)
Hence, PCA provides high-dimensional projection vectors preserving the original distance, similarity, or variance as much as possible. Although one can have the basis of the orthogonal projection using PCA, how can one interpret the resulting plot as a projection of hypercubes?
[…]
$\bm{e}_{i}=2\begin{bmatrix}u_{i;j}\\ u_{i;k}\end{bmatrix}.$ (10)
From now on, we call $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$ biplot bases. Notice that, when the data points are binary vertices $\bm{b}\in\left\{1,0\right\}^{N}$ of a system, the biplot bases exactly match the basis of the orthogonal projection of a hypercube, while when the data points are Ising vertices $\bm{s}\in\left\{+1,-1\right\}^{N}$, the biplot bases with doubled magnitudes exactly match the contribution bases. This difference arises from the difference in the length of hypercubic edges: they are one for binary vertices and two for Ising vertices. This is why we add factor $2$ in Eq. (10). By performing PCA and plotting biplot bases, one can obtain a reproducible, interpretable, and automatic projection of a hypercube.
IV Hypercubic PCA
In the previous Sec. III, we introduced PCA as a method to project hypercubes and presented the interpretation of the projection using biplot bases. PCA, in practice, requires the probability distribution $p\left(\bm{s}\right)$ of the vertices $\bm{s}$ to calculate the covariance matrix. How does $p\left(\bm{s}\right)$ affect the projection of hypercubes? How do the resulting coordinates of the vertices change when $p\left(\bm{s}\right)$ is varied? In this section, we project hypercubes using PCA while varying the probability distribution $p\left(\bm{s}\right)$ of the vertices. Throughout the examples, we reveal the trends of the resulting projections: the leading PC corresponds to the vertices with higher probability, and PCA distorts the distance between the vertices around the origin of the projected space.
Figure 4: Orthogonal projections of four-dimensional hypercubic vertices using PCA. (a) A projection of a four-dimensional hypercube where vertices are weighted randomly. Red filled circles are the vertices and lines are the edges of the hypercube. The magnitude of weight is proportional to the area of the vertex. Arrows are biplot bases and the boxes on the bottom right indicate the correspondence between the colors of the arrows and the original dimensions. (b) Fraction of explained variance by each PC of (a). (c) PC1 loading, and (d) PC2 loading of random weighted hypercubic vertices of (a). (e) Hamiltonian path on a four-dimensional hypercube in (a). For a different realization of random weight, see the Supplemental Material [105]. (f–j) Same as (a–e) but with bipolar distribution. Two of the vertices, $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$, are more weighted than the others.
IV.1 Random distribution
To begin with, we consider the random probability distribution of the vertices. In Fig. 4(a), we perform PCA on a four-dimensional hypercube with randomly weighted vertices. The probability distribution of the vertices is drawn randomly from a uniform distribution. The vertices with higher probabilities, which contribute more to the variance than the others, tend to lie on the outer part of the projection. Nevertheless, some less weighted vertices also lie on the outer part of the projection. We find that these vertices have counterparts with larger weights across the origin of the projected space. For example, in Fig. 4(a), the vertex around $\begin{bmatrix}0.4&-1.7\end{bmatrix}^{\top}$ has a small weight but lies on the outer part of the projection. Across the origin (around $\begin{bmatrix}-0.4&1.7\end{bmatrix}^{\top}$), it has a vertex with a large weight. Because PCA tries to preserve the variance of the data, the heavily weighted vertices tend to lie on the outer part of the projection, and sometimes less weighted vertices are also projected to the outer part of the projection.
The biplot bases in Fig. 4(a) are drawn as arrows from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$, assisting us in estimating the original binary coordinates of the vertices. Here, we abbreviate $+1$ to $+$ and $-1$ to --. For example, the projected vertex around $\begin{bmatrix}-2.0&-0.25\end{bmatrix}^{\top}$ corresponds to the vertex $\begin{bmatrix}-&-&+&-\end{bmatrix}^{\top}$ and the projected vertex around $\begin{bmatrix}-0.4&1.7\end{bmatrix}^{\top}$ corresponds to the vertex $\begin{bmatrix}+&-&+&+\end{bmatrix}^{\top}$. Equivalently, one can follow the hypercubic edge from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ (the origin of the biplot bases) to the vertex of interest by changing the corresponding elements indicated by the biplot bases to know the original Ising coordinate. Notice that it is not necessary to start from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ to infer the original Ising coordinate: if one knows the original coordinates of any vertex, one can infer the rest from the biplot bases, cf. Eq. (2).
We then examine the fraction of explained variance and loading of PCA in Fig. 4(a). We show the fraction of explained variance by each PC in Fig. 4(b). PC1 explains approximately 30% of the variance, followed by PC2 and PC3, which explain around 22% each, and PC4 explains just above 20%. Therefore, less than 60% of the variance is explained by the first two PCs. With a randomly weighted hypercube, the projection by the first two PCs does not explain a large fraction of the variance.
To understand the projection of the hypercube using PCA, we show the loading of PC1 and PC2 in Figs. 4(c) and 4(d). The PC1 loading [Fig. 4(c)] shows that variable 3 contributes negatively, but the rest contribute positively, and their absolute values are almost the same. We see that the sign of the PC1 loading corresponds to the vertices with the highest PC1 score, $\operatorname{sgn}\left(\bm{u}_{1}\right)=\begin{bmatrix}+&+&-&+\end{bmatrix}^{\top}$, indicating that the weighted vertices correspond to the PC1 loading. Here, $\operatorname{sgn}$ is the element-wise sign function. Unlike the PC1 loading, the PC2 loading [Fig. 4(d)] shows a different pattern of contribution for each variable. Although, similar to the PC1 loading, the sign of the PC2 loading corresponds to the vertices with the highest PC2 score, $\operatorname{sgn}\left(\bm{u}_{2}\right)=\begin{bmatrix}+&-&+&+\end{bmatrix}^{\top}$. Leading PC captures the weighted vertices, and the rest of the PCs capture the vertices with lower weights.
We then present the usage of the projection of the hypercube using PCA. In Fig. 4(e), the Hamiltonian path on the hypercube is shown following the biplot. Starting from the vertex $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$, one can follow the path to the vertex $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$ in the original high-dimensional space, by knowing which digit changed by following the biplot bases. These properties may be useful for understanding pathways of state transition of the target system. Notice, however, that the arrows indicating the Hamiltonian path overlap around the origin of the projected space. This overlap makes it difficult to follow the path. We address this issue in Sec. VI.
Figure 5: Orthogonal projections of four-dimensional hypercubic vertices using PCA. $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$ are the most weighted vertices, $\begin{bmatrix}-&-&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&-&-\end{bmatrix}^{\top}$ are the second most weighted vertices, $\begin{bmatrix}-&+&-&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&-&+&-\end{bmatrix}^{\top}$ are the third most weighted vertices, and the rest of them are weighted randomly. (a) Projection of the hypercube by PC1 and PC2. Red filled circles are the vertices and lines are the edges of the hypercube. The magnitude of weight is proportional to the area of the vertices. Arrows are biplot bases and the boxes on the bottom right show the correspondence between the colors of the arrows and the original dimensions. Dashed arrows correspond to the projection of original high-dimensional vectors, $\begin{bmatrix}-&+&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$. The original coordinate is indicated around the lower right of the arrowhead as an array of filled $\blacksquare$ (indicates $+$) or empty $\square$ (indicates --) boxes. Horizontal and vertical dashed lines crossing the origin are for visual aid. (b) Same as (a) but by PC1 and PC3. (c) Fraction of explained variance by each PC. (d) PC1 loading, (e) PC2 loading, and (f) PC3 loading.
IV.2 Bipolar distribution
Considering the results of PCA on randomly weighted hypercubic vertices, which indicate that weighted pairs of vertices play an important role in the projection, we perform PCA on hypercubic vertices with a bipolar weight distribution. In Fig. 4(f), we show the result of PCA on a four-dimensional hypercube where all vertices are weighted equally except for two of them. Two of the vertices $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$, which are the most distant from each other in the original four-dimensional space, are weighted more than the others. These weighted vertices are projected to have larger magnitudes of PC1 scores, but the rest of the vertices are projected in the order of Hamming distance from the weighted vertices along PC1: the resulting projection is the Hamming projection, cf. Fig. 2(c) and 2(d).
We show the fraction of explained variance by each PC in Fig. 4(g) to validate the projection of Fig. 4(f). Similar to the random weighted case in Fig. 4(b), PC1 explains approximately 30% of the variance. Unlike the random weighted case in Fig. 4(b), PC2 to PC4 have the same explained variance because of the uniform probability except for two vertices. The first two PCs explain a comparable proportion of variance (approximately 50%) to Fig. 4(b).
We arrive at the PC loadings of PC1 and PC2 in Figs. 4(h) and 4(i), where the former is expected to correspond to the weighted vertices. The PC1 loading in Fig. 4(h) shows that all variables contribute equally, contrary to the randomly weighted case in Fig. 4(c). This uniform PC1 loading in Fig. 4(h) supports that Fig. 4(f) is the Hamming projection. As expected, the most weighted vertices are projected to have the largest magnitude of PC1 scores, and the element-wise sign of the PC1 loading corresponds to the most weighted vertices, $\operatorname{sgn}\left(\bm{u}_{1}\right)=\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$. Similar to the randomly weighted PCA in previous Sec. IV.1, the element-wise sign of the PC2 loading in Fig. 4(i) corresponds to the vertices with the highest PC2 scores, $\begin{bmatrix}+&+&-&-\end{bmatrix}^{\top}$. While the PC1 loading relates to the weighted vertex pair, the PC2 loading seems to be randomly chosen due to the uniform probability distribution.
The projection of the Hamiltonian path shown in Fig. 4(j) is an example usage of the resulting Hamming projection. In addition to the traceability—visualized original dimension—of the Hamiltonian path as in Fig. 4(e), one can see how the transition on the hypercubic edge relates to the Hamming distance from the vertices $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$.
IV.3 Sexapolar distribution
Motivated by the results of PCA on bipolar weighted hypercubic vertices, we expect that the projection of the hypercube by PCA reflects the weighted vertices. In Fig. 5, we perform PCA on a four-dimensional hypercube where three pairs of the most distant vertices are weighted more, and the rest are weighted randomly. Each of the three pairs of vertices is weighted differently so that the most weighted vertices are $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$, the second most weighted vertices are $\begin{bmatrix}-&-&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&-&-\end{bmatrix}^{\top}$, and the third most weighted vertices are $\begin{bmatrix}-&+&-&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&-&+&-\end{bmatrix}^{\top}$. These weighted pairs are selected to be perpendicular to each other.
We show the projection by the first two PCs in Fig. 5(a), finding a similar trend as the bipolar weighted PCA in previous Sec. IV.2. The most weighted vertex pair is projected to have a larger magnitude of PC1 score as in Fig. 4(f). The second most weighted vertex pair is projected to have a larger magnitude of PC2 score. Notice that the third most weighted vertex pair is projected around the origin, one is around $\begin{bmatrix}-0.1&0.25\end{bmatrix}^{\top}$ and the other is around $\begin{bmatrix}0.1&-0.25\end{bmatrix}^{\top}$, despite being the most distant from each other in the original space. This can be understood by the minimum inner-product-error formulation of PCA. In this formulation, each hypercubic vertex is projected so that the inner product between the vertices is preserved. Thus, a vertex should be projected far from the most distant vertex across the origin. That is why the most weighted vertex pair is projected to have a larger magnitude of the PC1 (or PC2 for the second weighted vertex pair) score. Then the two PCs are already used to locate the two most weighted vertex pairs, and the third most weighted vertices, which are perpendicular to the most and second most weighted vertices in our example, are both projected around the origin even though they are far from each other in the original high-dimensional space. Moreover, neighbors in the original high-dimensional space are projected to be distant from each other in the projected space. For instance, we emphasize the two vertices, $\begin{bmatrix}-&+&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$, in Fig. 5(a). These vertices are projected to be far from each other in the projected space though they are neighboring vertices in the original space. Similarly, the vertices $\begin{bmatrix}-&+&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&-&+&+\end{bmatrix}^{\top}$ are located close to each other around $\begin{bmatrix}1&1\end{bmatrix}^{\top}$, but they are not neighboring vertices in the original space (the Hamming distance between them is two). These results show that the neighboring relationships in the original space are not necessarily preserved in the projected space: the distances between the vertices in the projected space can be misleading. This is the reason why distance-preserving formulation of PCA, i.e., the idea behind classical MDS, is not suitable to interpret the projections of hypercubes.
When we plot the projection by PC1 and PC3 in Fig. 5(b) instead of by PC1 and PC2, we see that the third most weighted vertex pair is projected to have a larger magnitude of PC3 score but the second most weighted vertex pair is projected around the origin. Our results indicate that each PC loading represents the weighted vertex pair and the PC score corresponds to the similarity to or distance from the weighted vertex pair.
We validate the projection of Figs. 5(a) and 5(b) by the explained variance of each PC. We show, in Fig. 5(c), the fraction of explained variance by each PC. The fraction gradually decreases as the PC number increases, unlike the random weighted case [Fig. 4(b)] and the bipolar weighted case [Fig. 4(g)]. The explained variance by the first two PCs (more than 60%) slightly increases more than that of the random weighted case [Fig. 4(b)] and the bipolar weighted case [Fig. 4(g)].
We show the loading of PC1, PC2, and PC3 in Figs. 5(d), 5(e), and 5(f), respectively, to examine how the PCs represent the weighted vertices. Similar to the bipolar weighted case, the element-wise sign of PC1 loading corresponds to the most weighted vertices: $\operatorname{sgn}\left(\bm{u}_{1}\right)=\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$. In addition, the element-wise sign of PC2 (PC3) loading also corresponds to the second (third) most weighted vertices. Due to the noise from the randomly weighted vertices, PC1 loading is not completely uniform, and the projection is not exactly the Hamming projection, but we find a similar trend.
IV.4 Brief review and preview
In all orthogonal projections [Figs. 4(a), 4(f), 5(a), and 5(b)] of hypercubes using PCA, the weighted vertices have high PC1 scores, indicating that PC1 is associated with the weighted vertices. This observation is discussed further in Sec. V.
We have observed that some neighboring relationships between the vertices are distorted: the distances between the vertices in the projected space can be misleading. When the weighted vertex pairs are perpendicular to each other [Fig. 5(a) and 5(b)], the second most weighted vertices are projected to have high PC2 scores, and the third most weighted vertices are projected to have high PC3 scores. We discuss this further in Sec. VI.1 from the perspective of projection quality and the inner-product error.
When two PC loadings cannot capture all the weighted vertices, the missed weighted vertices can be placed near the origin of the projected space. This centrality of the projected vertices is discussed further in Sec. VI.2 and its effect on the inner-product error is discussed in Sec. VI.3.
V PC1 loading and weighted vertices
In Sec. IV, we find that the element-wise sign of the PC1 loading corresponds to the vertices with the highest PC scores—the weighted vertices. The results indicate that the PC1 loading is related to the weighted vertices, and the PC1 score corresponds to the similarity to or distance from the weighted vertices. To understand this, in this section, we analytically and numerically examine the properties of PCA, particularly the correspondence between the leading PC loading and the weighted vertices. We first consider the ideal probability distribution, then expand our consideration to more general distributions. We also numerically validate the analytical results.
V.1 Analytical investigation
We consider an ideal situation where the distribution is concentrated on a few hypercubic vertices. Consider, for example, a low-temperature canonical ensemble for an Ising spin system with $N$ spins, where the vertex distribution is dominated by a weighted vertex (or ground state) $\bm{\xi}\in\left\{+1,-1\right\}^{N}$ and its globally spin-flipped vertex (state) $-\bm{\xi}$. Suppose the distribution is idealized as a bipolar one
$p\left(\bm{s}\right)\approx\frac{1}{2}\left(\delta_{+\bm{\xi},\bm{s}}+\delta_{-\bm{\xi},\bm{s}}\right),$ (11)
where $\delta_{\bm{y},\bm{x}}$ is the Kronecker delta function for vectors, which is equal to $1$ when $\bm{x}=\bm{y}$ and 00 otherwise. Given that the mean vector is the zero vector, $\bm{s}\right>=\sum_{\bm{s}}\frac{1}{2}\left(\delta_{+\bm{\xi},\bm{s}}+\delta_{-\bm{\xi},\bm{s}}\right)\bm{s}=\bm{0},$ the covariance matrix [Eq. (7)] becomes999 If the distribution is unipolar, $p\left(\bm{s}\right)\approx\delta_{+\bm{\xi},\bm{s}},$ the covariance matrix becomes $\bm{\varSigma}=\sum_{\bm{s}}\delta_{+\bm{\xi},\bm{s}}\left(\bm{s}-\bm{\xi}\right)\left(\bm{s}-\bm{\xi}\right)^{\top}=\bm{0}\bm{0}^{\top},$ which results in all vertices being projected to the origin.
$\bm{\varSigma}=\sum_{\bm{s}}\frac{1}{2}\left(\delta_{+\bm{\xi},\bm{s}}+\delta_{-\bm{\xi},\bm{s}}\right)\bm{s}\bm{s}^{\top}=\bm{\xi}\bm{\xi}^{\top},$ (12)
which is already diagonalized with the eigenvector $\bm{\xi}$ and the corresponding eigenvalue $N$, i.e., $\bm{\varSigma}\bm{\xi}=\bm{\xi}\bm{\xi}^{\top}\bm{\xi}=N\bm{\xi}$, where $N$ is the largest eigenvalue. Therefore, the PC1 loading is
$\bm{u}_{1}=\frac{1}{\sqrt{N}}\bm{\xi},$ (13)
which satisfies $\bm{\varSigma}\bm{u}_{1}=N\bm{u}_{1}$. Notice the normalization factor $\frac{1}{\sqrt{N}}=\frac{1}{\sqrt{\bm{\xi}^{\top}\bm{\xi}}}$. Also, the element-wise sign of the PC1 loading is the weighted vertex vector, $\operatorname{sgn}\left(\bm{u}_{1}\right)=\bm{\xi}$. This explains why we observe that the element-wise sign of the PC loading is the same as the weighted state. If we remove the normalization factor, the PC loading is exactly same as the weighted state.
The projected coordinates on PC1 (PC1 score) become
$r_{1}\left(\bm{s}\right)=\bm{u}_{1}^{\top}\bm{s}=\frac{1}{\sqrt{N}}\bm{\xi}^{\top}\bm{s}=\sqrt{N}Q\left(\bm{\xi},\bm{s}\right),$ (14)
[…]
$D_{\mathrm{H}}\left(\bm{b},\bm{b}^{\prime}\right)\coloneqq\sum_{i=1}^{N}\left(1-\delta_{b_{i},b_{i}^{\prime}}\right).$ (15)
Here, $\delta_{b_{i},b_{i}^{\prime}}$ is the Kronecker delta function for scalars. Because $s_{i}s^{\prime}_{i}\in\left\{+1,-1\right\}$ and $\delta_{b_{i},b_{i}^{\prime}}=\frac{1}{2}\left(1+s_{i}s^{\prime}_{i}\right)$, the Hamming distance $D_{\mathrm{H}}\left(\bm{s},\bm{s}^{\prime}\right)$ between two Ising state vectors $\bm{s}$ and $\bm{s}^{\prime}$ is
$D_{\mathrm{H}}\left(\bm{s},\bm{s}^{\prime}\right)=$ $\sum_{i=1}^{N}\left(1-\frac{1+s_{i}s^{\prime}_{i}}{2}\right)$
$=$ $\frac{N-\bm{s}^{\top}\bm{s}^{\prime}}{2}=N\frac{1-Q\left(\bm{s},\bm{s}^{\prime}\right)}{2}.$ (16)
Using this Eq. (16), we have $Q\left(\bm{\xi},\bm{s}\right)=\frac{N-2D_{\mathrm{H}}\left(\bm{\xi},\bm{s}\right)}{N}$. Then, the PC1 score of Eq. (14) is written as
$r_{1}\left(\bm{s}\right)=\frac{+N-2D_{\mathrm{H}}\left(+\bm{\xi},\bm{s}\right)}{\sqrt{N}}$ (17)
or using $r_{1}\left(\bm{s}\right)=-\left(-\frac{1}{\sqrt{N}}\bm{\xi}\right)\bm{s}=-\sqrt{N}Q\left(-\bm{\xi},\bm{s}\right)$ and $Q\left(-\bm{\xi},\bm{s}\right)=\frac{N-2D_{\mathrm{H}}\left(-\bm{\xi},\bm{s}\right)}{N}$,
$r_{1}\left(\bm{s}\right)=\frac{-N+2D_{\mathrm{H}}\left(-\bm{\xi},\bm{s}\right)}{\sqrt{N}}.$ (18)
The PC1 score of any sample state $\bm{s}$ is then linearly equivalent to the Hamming distance between a state $\bm{s}$ and the weighted state $\bm{\xi}$ (or $-\bm{\xi}$). Therefore, the distribution of Eq. (11) guarantees the Hamming projection.
We extend our consideration to a more general distribution by perturbation theory of quantum mechanics. When the distribution is approximately quadripolar, i.e.,
$p\left(\bm{s}\right)\approx\frac{1-\epsilon}{2}\left(\delta_{+\bm{\xi}_{1},\bm{s}}+\delta_{-\bm{\xi}_{1},\bm{s}}\right)+\frac{\epsilon}{2}\left(\delta_{+\bm{\xi}_{2},\bm{s}}+\delta_{-\bm{\xi}_{2},\bm{s}}\right),$ (19)
with perturbation parameter $0\leq\epsilon\ll 1$. Using perturbation theory, we obtain the PC1 loading as
$\bm{u}_{1}\propto\bm{\xi}_{1}$ (20)
and PC2 loading as
$\bm{u}_{2}\propto\bm{\xi}_{2}-\frac{1}{N}\bm{\xi}_{1}^{\top}\bm{\xi}_{2}\bm{\xi}_{1}.$ (21)
When the two weighted vertices are perpendicular to each other, $\bm{\xi}_{1}^{\top}\bm{\xi}_{2}=0$, the PC2 score linearly depends on the Hamming distance from the second most weighted vertices, similar to the PC1 score. The derivation is in Appendix C.
V.2 Numerical validation
With the ideal distribution of Eq. (11), we show that the PC loading corresponds to the weighted vertices. This argument is expected to apply to distributions approximately similar to Eq. (11), but how valid is this assumption?
[…]
$\Delta u_{j;i}\coloneqq\sqrt{\frac{1}{N}\sum_{j=1}^{N}\left({u_{j;i}}-\frac{1}{N}\sum_{j=1}^{N}u_{j;i}\right)^{2}},$ (22)
where $u_{j;i}$ is the $j$th element of the $i$th PC loading. If all elements of PC loading are the same, the standard deviation is zero.
In Fig. 6, we show the standard deviation dependency on the weight parameter $z$. At $z=\frac{1}{2^{N}}$, where all vertices have the same weight, the standard deviation of PC1 loading abruptly changes to zero, indicating that PC1 is aligned with the selected state $\bm{\xi}$ (or $-\bm{\xi}$). This indicates that as long as the distribution of the vertices is qualitatively similar to Eq. (11), the PC1 is aligned with the selected state $\bm{\xi}$ (or $-\bm{\xi}$). We find that the last PC also shows a sudden change in the standard deviation at $z=\frac{1}{2^{N}}$. The opposite happens to the last PC; the last PC is aligned with the selected state when the selected state is weighted less than the others. Thus, the PC1 is robustly equivalent to the weighted vertices as long as they are weighted more than the others.
Figure 6: The standard deviation of PC loading $\Delta u_{j;i}$ dependency on the weight parameter $z$. We performed PCA on the vertices of the six-dimensional hypercube varying the weight $z$ of the selected state $\bm{\xi}$. The selected vertices are $\bm{\xi}=\begin{bmatrix}+&+&+&+&+&+\end{bmatrix}^{\top}$ and $-\bm{\xi}=\begin{bmatrix}-&-&-&-&-&-\end{bmatrix}^{\top}$. Vertical dashed lines indicate $z=\frac{1}{2^{6}}$, where all vertices have the same weight. For clarity of visualization, we plot a limited number of sampled points of the weight parameter $z$, but we also sampled more densely and confirmed the same trend.
Figure 7: Quality of various projections of hypercubes. The inner product in the original space $\bm{s}^{\top}\bm{s}^{\prime}$ is plotted as a function of the inner product in the projected space $\bm{r}^{\top}\left(\bm{s}\right)\bm{r}\left(\bm{s}^{\prime}\right)$. (a) The isometric projection of a cube in Fig. 2(a). (b) The isometric projection of a tesseract in Fig. 2(b). (c) The Hamming projection of a cube in Fig. 2(c). (d) The Hamming projection of a tesseract in Fig. 2(d). (e) The fractal projection of a six-dimensional hypercube in Fig. 2(e). (f) The fractal projection of an eight-dimensional hypercube in Fig. 2(f). (g) The projection by PC1 and PC2 of a randomly weighted tesseract in Fig. 4(a). (h) The projection by PC1 and PC2 of a bipolarly weighted tesseract in Fig. 4(f). (i) The projection by PC1 and PC2 of a sexapolarly weighted tesseract in Fig. 5(a). (j) The projection by PC1 and PC3 of a sexapolarly weighted tesseract in Fig. 5(b). The dashed line in each panel crosses the origin with slope one. In (g–j), we show the squared sum of unexplained variance as a shade for the confidence interval. The width of the shade is $2\sum_{i\notin\left\{1,2\right\}}\lambda_{i}$ for (g–i), and $2\sum_{i\notin\left\{1,3\right\}}\lambda_{i}$ for (j). The root mean squared inner-product error is shown at the bottom right. Note that before the calculation of the inner product, we centered all variables to the origin, i.e., $\bm{s}\leftarrow\bm{s}-\left<\bm{s}\right>$ and $\bm{r}\left(\bm{s}\right)\leftarrow\bm{r}\left(\bm{s}\right)-\left<\bm{r}\left(\bm{s}\right)\right>$.
VI Dependency of inner-product error on projected coordinates and its inevitability
We have introduced several methods to project hypercubes and shown that PCA has several advantages. The remaining question is the quality of these projections. In this section, we compare the quality of the projections we have introduced so far. Through the investigation of quality, we observe that the error of the projection arises from the vertices located around the center of the projected space. We then theoretically explain the reason for this tendency by examining the number of vertices along PC1 and the upper bound of the inner-product error.
VI.1 Quality of the projections
To evaluate the quality of orthogonal projections of hypercubes, we investigate the inner-product error [Eq. (5)] between the original space and the projected space for all possible pairs of vertices of hypercubes. As mentioned in Sec. III.1, inner-product error is minimized in PCA. It indicates which pairs of vertices are responsible for errors. Inner-product error can also measure the quality of the projection in general because it indicates how the similarity between original and projected vertices is preserved.
In Fig. 7, we show the inner products of all possible pairs of vertices in the original space as a function of those in the projected space. In general, the difference between the inner product in the original space and those in the projected space becomes larger as the inner product in the projected space approaches zero, even if the projection method is not PCA. Exceptions are found in the fractal projections [Figs. 2(e) and 2(f)], where the inner products in the original space and projected space do not match even if the inner product in the projected space is relatively larger. All the projections show a similar trend that the inner-product error and the error created by fractal projections is larger than the others. The inner product becomes zero in two conditions: when two vectors are orthogonal, or when one or both of them are zero vectors. As Fig. 5(b) shows, the latter contributes more than the former to the inner-product error because if both vertices of a pair have large probability, the inner-product error becomes large. In fact, the inner-product error increases (cf. inset of Figs. 7(i) and 7(j)) when the weighted vertices are projected around the origin [Figs. 5(b)] compared to when they are projected far from the origin [Figs. 5(a) and 7(i)].
Because of this general trend of the inner-product error, orthogonal projections of hypercubes can be misleading for the vertices located around the center of the projected space. The pairwise inner product between the vertices both located around the center of the projected space might be the most distant pair in the original space. For example, as we mentioned earlier, regarding Fig. 5 in Sec. IV, the orthogonal projection of a hypercube using PCA locates the third (or second) weighted vertex pair around the origin, but their original Ising coordinates are the most distant pair. In the same way, inspecting other types of orthogonal projections in Figs. 2 and 3, we find that the vertices located and overlapped around the center of the projected space can be the most distant in the original space. Why is this trend ubiquitous in the orthogonal projection of hypercubes? We answer this question in Sec. VI.2 and Sec. VI.3.
VI.2 Centrality of the projection
Several projections of hypercubes, such as Figs. 2(a), 2(d), 2(e), 3(a), 4(a), 4(f), 5(a), and 5(b), show that multiple vertices are located around the origin of the projected space. In Fig. 7, we find that the pairs of vertices including those around the origin contribute to the inner-product error. It seems ubiquitous that a number of vertices projected around the origin in the orthogonal projection of hypercubes is significant. Why are some vertices projected to accumulate around the origin? To answer this question, we consider the number of vertices along the horizontal axis of the Hamming projection. We show that at large $N$, the normalized number of vertices having Hamming distance $D_{\mathrm{H}}$ is approximated by the zero-mean Gaussian (normal) distribution with appropriate centering. In other words, we show that the distribution of the PC1 score of unweighted hypercubic vertices using the PC1 loading of Eq. (13) is the Gaussian distribution.
We first consider the number of vertices along the Hamming-distance axis of the Hamming projection. The projected coordinates with Hamming distance $D_{\mathrm{H}}=D_{\mathrm{H}}\left(-\bm{\xi},\bm{s}\right)$ are given by Eq. (18), $r\left(D_{\mathrm{H}}\right)\in\left\{-\sqrt{N}+\frac{2}{\sqrt{N}}D_{\mathrm{H}}\right\}_{D_{\mathrm{H}}=0}^{N}$. Because the Hamming distance is the number of unmatched elements in the binary vectors, the binomial coefficient $\binom{N}{D_{\mathrm{H}}}\coloneqq\frac{N!}{D_{\mathrm{H}}!\left(N-D_{\mathrm{H}}\right)!}$ gives the number of vertices at each possible $r\left(D_{\mathrm{H}}\right)$. The normalized number of vertices with the Hamming distance $D_{\mathrm{H}}$ is $\varrho\left(D_{\mathrm{H}}\right)=\binom{N}{D_{\mathrm{H}}}\frac{1}{2^{N}}=\binom{N}{D_{\mathrm{H}}}\left(\frac{1}{2}\right)^{D_{\mathrm{H}}}\left(1-\frac{1}{2}\right)^{N-D_{\mathrm{H}}}$, i.e., the binomial distribution. With large $N$, $\varrho\left(D_{\mathrm{H}}\right)$ is asymptotically a Gaussian distribution with mean $\frac{N}{2}$ and variance $\frac{N}{4}$,
$\varrho\left(D_{\mathrm{H}}\right)\sim\frac{1}{\sqrt{2\uppi\frac{N}{4}}}\exp\left[-\frac{1}{2}\frac{\left(D_{\mathrm{H}}-\frac{N}{2}\right)^{2}}{\frac{N}{4}}\right].$ (23)
By changing the variable using Eq. (18), the distribution as a function of the projected coordinate $r$, $\varrho\left(r\right)$, is obtained:
$\varrho\left(r\right)=\frac{\mathrm{d}D_{\mathrm{H}}}{\mathrm{d}r}\varrho\left(D_{\mathrm{H}}\right)\simeq\frac{1}{\sqrt{2\uppi}}\exp\left(-\frac{1}{2}r^{2}\right),$ (24)
which is the standard Gaussian distribution. Thus, as the dimension of the hypercube $N$ increases, a larger number of vertices are projected around the origin in the Hamming projection.
For the Hamming projection, we find that the distribution of projected coordinates follows the Gaussian distribution. To what extent is this result valid? Our numerical results [Fig. 6] in Sec. V.2 indicate that as long as the distribution of the vertices is qualitatively similar to Eq. (11), the Hamming projection is guaranteed. Thus, the number of vertices along the Hamming-distance axis (PC1 loading) is approximated by a distribution close to the Gaussian distribution if the distribution of the vertices has bipolarity. This is the reason why the number of vertices along PC1 loading tends to follow the Gaussian distribution even if the distribution is not ideally bipolar [78].
In general, the distribution of hypercubic vertices along a linear projection axis can be shown to be Gaussian under fairly weak assumptions. First, note that any transformation vector $\bm{v}$ can be represented as the normalization $\bm{v}=\frac{\bm{w}}{\left|\bm{w}\right|}$ of a weighted superposition of non-overlapping binary sub-states,
$\bm{w}=\sum_{g=1}^{M_{\mathrm{G}}}\frac{a_{g}}{\sqrt{n_{g}}}\bm{\xi}_{g}.$ (25)
Here, $M_{\mathrm{G}}$ is the number of sub-states, reflecting the complexity of the original transformation vector, and $\bm{\xi}_{g}\in\left\{+1,-1,0\right\}^{N}$ is the $g$th sub-state with $n_{g}$ non-zero elements and a weight $a_{g}$. Note that, as $\left|\bm{\xi}_{g}\right|=\sqrt{n_{g}}$ by definition and the sub-states do not overlap with each other, $\bm{\xi}_{g}^{\top}\bm{\xi}_{g^{\prime}}=n_{g}\delta_{g,g^{\prime}}$, the squared norm of $\bm{w}$ is the sum of the squared weights: $\left|\bm{w}\right|^{2}=\sum_{g=1}^{M_{\mathrm{G}}}a_{g}^{2}$. The projected coordinate for a vertex $\bm{s}$ then is expressed as a weighted superposition of sub-coordinates,
$r=\frac{1}{\left|\bm{w}\right|}\bm{w}^{\top}\bm{s}=\frac{1}{\left|\bm{w}\right|}\sum_{g=1}^{M_{\mathrm{G}}}a_{g}r_{g},$ (26)
where $r_{g}$ is the contribution from the $g$th sub-state,
$r_{g}=\frac{1}{\sqrt{n_{g}}}\bm{\xi}_{g}^{\top}\bm{s}=\frac{-n_{g}+2D_{\mathrm{H}}^{\left(g\right)}}{\sqrt{n_{g}}}.$ (27)
$D_{\mathrm{H}}^{\left(g\right)}$ is the Hamming distance between the non-zero dimensions of the $g$th sub-state $-\bm{\xi}_{g}$ and the corresponding dimensions of a vertex $\bm{s}$, as in Eq. (18). If the non-zero dimension $n_{g}$ of all sub-states is sufficiently large, the distribution of the projected sub-coordinate $r_{g}$ asymptotically becomes the standard Gaussian with the same procedure leading to Eqs. (23) and (24). Therefore, the projected coordinate $r$, which is the weighted sum of the sub-coordinates, is also Gaussian with zero mean, and its variance is unity:
$r^{2}\right>_{\varrho\left(r\right)}=\frac{1}{\left|\bm{w}\right|^{2}}\sum_{g=1}^{M_{\mathrm{G}}}a_{g}^{2}\left<r_{g}^{2}\right>_{\varrho\left(r_{g}\right)}=1.$ (28)
VI.3 Inner-product error bounds of projections
We then evaluate the dependency of the inner-product error on the projected coordinates. Suppose the hypercubic vertices are projected to form a Hamming projection, and the projected coordinates are given as in Eq. (17). The Hamming distance between the two vertices $\bm{s}$ and $\bm{s}^{\prime}$ satisfies the triangular inequalities:
$D_{\mathrm{H}}\left(\bm{s},\bm{s}^{\prime}\right)\leq$ $D_{\mathrm{H}}\left(+\bm{\xi},\bm{s}\right)+D_{\mathrm{H}}\left(+\bm{\xi},\bm{s}^{\prime}\right)$ (29)
$=$ $N-\frac{\sqrt{N}}{2}\left[r_{1}\left(\bm{s}\right)+r_{1}\left(\bm{s}^{\prime}\right)\right],$ (30)
and
$D_{\mathrm{H}}\left(\bm{s},\bm{s}^{\prime}\right)\leq$ $D_{\mathrm{H}}\left(-\bm{\xi},\bm{s}\right)+D_{\mathrm{H}}\left(-\bm{\xi},\bm{s}^{\prime}\right)$ (31)
$=$ $N+\frac{\sqrt{N}}{2}\left[r_{1}\left(\bm{s}\right)+r_{1}\left(\bm{s}^{\prime}\right)\right].$ (32)
Notice that from Eqs. (17) and (18), $D_{\mathrm{H}}\left(\pm\bm{\xi},\bm{s}\right)=\frac{N\mp\sqrt{N}r_{1}\left(\bm{s}\right)}{2}$. Combining Eqs. (30) and (32), we obtain
$D_{\mathrm{H}}\left(\bm{s},\bm{s}^{\prime}\right)\leq N-\frac{\sqrt{N}}{2}\left|r_{1}\left(\bm{s}\right)+r_{1}\left(\bm{s}^{\prime}\right)\right|.$ (33)
Using Eqs. (16), (17), and (33), the inner-product error of Eq. (5) satisfies
$\varepsilon(\bm{s},\bm{s}^{\prime})=$ $N-2D_{\mathrm{H}}\left(\bm{s},\bm{s}^{\prime}\right)-r_{1}\left(\bm{s}\right)r_{1}\left(\bm{s}^{\prime}\right)$
$\geq$ $-N+\sqrt{N}\left|r_{1}\left(\bm{s}\right)+r_{1}\left(\bm{s}^{\prime}\right)\right|-r_{1}\left(\bm{s}\right)r_{1}\left(\bm{s}^{\prime}\right),$ (34)
which is the lower bound of the inner-product error. Because the projected coordinate has linearity under the reflection of $\bm{s}$, i.e., $r_{1}\left(-\bm{s}\right)=-r_{1}\left(\bm{s}\right)$, the inner-product error of Eq. (5) has bilinearity, $\varepsilon\left(-\bm{s},\bm{s}^{\prime}\right)=\varepsilon\left(\bm{s},-\bm{s}^{\prime}\right)=-\varepsilon\left(\bm{s},\bm{s}^{\prime}\right)$. Thus, from Eq. (34), $-\varepsilon\left(\bm{s},\bm{s}^{\prime}\right)=\varepsilon\left(\bm{s},-\bm{s}^{\prime}\right)\geq-N+\sqrt{N}\left|r_{1}\left(\bm{s}\right)+r_{1}\left(-\bm{s}^{\prime}\right)\right|+r_{1}\left(\bm{s}\right)r_{1}\left(-\bm{s}^{\prime}\right)$, which is equivalent to the upper bound of the inner-product error,
$\varepsilon(\bm{s},\bm{s}^{\prime})\leq+N-\sqrt{N}\left|r_{1}\left(\bm{s}\right)-r_{1}\left(\bm{s}^{\prime}\right)\right|-r_{1}\left(\bm{s}\right)r_{1}\left(\bm{s}^{\prime}\right).$ (35)
We investigate the error between the vertices which are projected to be overlapped. If $r_{1}\left(\bm{s}\right)=r_{1}\left(\bm{s}^{\prime}\right)=l$, the inner-product error satisfies
$-N+2\sqrt{N}\left|l\right|-l^{2}\leq\varepsilon\leq N-l^{2}.$ (36)
Thus, the squared inner-product error is then bounded by
$0\leq\varepsilon^{2}\leq\left(N-l^{2}\right)^{2}.$ (37)
We normalize Eq. (37) by $N^{2}$, resulting in
$0\leq\left(\frac{\varepsilon}{N}\right)^{2}\leq\left(1-\frac{l^{2}}{N}\right)^{2}.$ (38)
We plot the upper bound of Eq. (38) in Fig. 8(a). We indeed find that the error can deviate most from zero when the two vertices are projected around the origin $l=0$.
We extend our consideration to the two-dimensional projection by PC1 and PC2, with orthogonal weighted states $\bm{\xi}_{1}$ and $\bm{\xi}_{2}$. Assume the probability distribution of Eq. (19), suppose we have a Hamming projection which is projected by Eqs. (20) and (21). For each dimension $r_{i}$ of the projected coordinates [Eq. (14)], the inequality of Eq. (33) is satisfied. We then combine them as a single inequality,
$2D_{\mathrm{H}}\left(\bm{s},\bm{s}^{\prime}\right)$
$\leq$ $2N-\frac{\sqrt{N}}{2}\left[\left|r_{1}\left(\bm{s}\right)+r_{1}\left(\bm{s}^{\prime}\right)\right|+\left|r_{2}\left(\bm{s}\right)+r_{2}\left(\bm{s}^{\prime}\right)\right|\right]$
$=$ $2N-\frac{\sqrt{N}}{2}\left|\bm{r}\left(\bm{s}\right)+\bm{r}\left(\bm{s}^{\prime}\right)\right|_{1}.$ (39)
Using the same procedure to derive Eqs. (34) and (35), the inner-product error of Eq. (5) satisfies
$-N+\sqrt{N}\left|\bm{r}\left(\bm{s}\right)+\bm{r}\left(\bm{s}^{\prime}\right)\right|_{1}-\bm{r}^{\top}\left(\bm{s}\right)\bm{r}\left(\bm{s}^{\prime}\right)$
$\leq$ $\varepsilon\left(\bm{s},\bm{s}^{\prime}\right)$
$\leq$ $+N-\sqrt{N}\left|\bm{r}\left(\bm{s}\right)-\bm{r}\left(\bm{s}^{\prime}\right)\right|_{1}-\bm{r}^{\top}\left(\bm{s}\right)\bm{r}\left(\bm{s}^{\prime}\right).$ (40)
Our interest is the inner-product error between vertices sharing the same projected coordinates. If $\bm{r}\left(\bm{s}\right)=\bm{r}\left(\bm{s}^{\prime}\right)=\bm{l}\coloneqq\begin{bmatrix}l_{1}&l_{2}\end{bmatrix}^{\top}$, Eq. (40) becomes
$-N+2\sqrt{N}\left|\bm{l}\right|_{1}-\bm{l}^{\top}\bm{l}\leq\varepsilon\leq N-\bm{l}^{\top}\bm{l}.$ (41)
Thus, the squared inner-product error is bounded by
$0\leq\varepsilon^{2}\leq\left(N-\bm{l}^{\top}\bm{l}\right)^{2},$ (42)
which is the extension of Eq. (37) to the two-dimensional projection. We normalize Eq. (42) by $N^{2}$, obtaining
$0\leq\left(\frac{\varepsilon}{N}\right)^{2}\leq\left(1-\frac{\bm{l}^{\top}\bm{l}}{N}\right)^{2}.$ (43)
We plot Eq. (43) in Fig. 8(b) and find that the error can be the largest when the two vertices are projected around the origin $\bm{l}=\bm{0}$.
Figure 8: The upper bound of the squared inner-product error of Hamming projections. (a) The upper bound [Eq. (38)] of the squared inner-product error between the vertices sharing the same projected coordinate $l$. (b) The upper bound [Eq. (43)] of the squared inner-product error between the vertices sharing the same projected coordinates $\bm{l}$. The hypercubic vertices are projected inside the dashed circle [Eq. (44)] in general, and in case of Hamming projection, inside the dotted square [Eqs. (45) and (46)].
Notice that the corners of Fig. 8(b) have a slightly higher bound, but here we show that such regions cannot be used as projected coordinates. The possible range of $r_{1}\left(\bm{s}\right)$ and $r_{2}\left(\bm{s}\right)$ is $\left[-\sqrt{N},\sqrt{N}\right]$. In Appendix D, however, we show that projected coordinates are limited to a specific region of
$\left|\bm{r}\left(\bm{s}\right)\right|\leq\sqrt{N}.$ (44)
We plotted the boundary of the possible region with a white dashed line in Fig. 8(b). Intuitively, this corresponds to the fact that all hypercubic vertices are on the surface of the $N$-dimensional sphere with radius $\sqrt{N}$, and the projection of the particular slice of the sphere is the possible region. Furthermore, the Hamming projection with two perpendicular weighted states has a tighter bound on the possible region,
$\left|r_{1}\left(\bm{s}\right)+r_{2}\left(\bm{s}\right)\right|\leq\sqrt{N}$ (45)
and
$\left|r_{1}\left(\bm{s}\right)-r_{2}\left(\bm{s}\right)\right|\leq\sqrt{N},$ (46)
which is drawn with a white dotted line in Fig. 8(b). The square shape of the possible region is due to the fact that four vertices are chosen to form a square in the Hamming projection as an idealization. We derive Eqs. (45) and (46) in Appendix D. Thus, only a limited region of the $r_{1}\left(\bm{s}\right)$-$r_{2}\left(\bm{s}\right)$ plane in the Hamming projection can be the projected coordinates.
What we show here is the reason that many vertices are projected around the origin in the orthogonal projection—whether by PCA or not—of hypercubes. Upon linear projection, hypercubic vertices concentrate near the origin of the projected space in an approximate Gaussian shape. This is why the inner-product error gets a dominant contribution from the vertices located around the center of the projected space. Indeed, the upper bound of the inner-product error is the largest when the vertices are projected around the origin. Our discussion in Secs. VI.2 and VI.3 entirely relies on the assumption that the hypercubic vertices are projected to form a Hamming projection, which is a rough approximation and might not be the case in general. The full properties of such non-Hamming projections cannot be revealed analytically, but as we show in Sec. V.2, the qualitative aspects of such projections are expected to be similar to those of the Hamming projection. Thus, the quality of the projection of hypercubes is typically worse around the center of the projected space.
VII Applications
So far, we have examined the orthogonal projection of hypercubes by various methods with particular emphasis on PCA. Through our investigation, we have obtained insights into the projection using PCA, which enables us to interpret the resulting projections. Here, employing several Ising spin systems, we apply the orthogonal projection of hypercubes using PCA, aiming to obtain the physical interpretation of them.
VII.1 The finite artificial kagome spin-ice system
To demonstrate the usage of the orthogonal projection of hypercubes using PCA, we apply our method to statistical mechanical models. Specifically, we employ the hypercubic energy landscape [69] of finite artificial spin-ice systems [128]. For this purpose, we consider the paradigmatic Ising spin system on the kagome lattice because the kagome spin-ice systems can be mapped to the antiferromagnetic kagome Ising spin system [129]. In Fig. 9, we illustrate the finite artificial spin-ice systems and corresponding Ising spin interaction networks to be considered: one-ring [Figs. 9(a1) and 9(b1)], two-ring [Figs. 9(a2) and 9(b2)], and three-ring [Figs. 9(a3) and 9(b3)] systems. These systems were experimentally realized [128, 69], and the observed dynamics were analyzed through hypercubic energy landscapes [69] but not with PCA.
[…]
The Hamiltonian of the Ising spin system is defined as $\mathcal{H}\left(\bm{s}\right)\coloneqq-\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}s_{i}J_{i,j}s_{j}-\sum_{i=1}^{N}s_{i}h_{i}=-\frac{1}{2}\bm{s}^{\top}\bm{J}\bm{s}-\bm{s}^{\top}\bm{h},$ where the state vector is defined as $\bm{s}\coloneqq\begin{bmatrix}s_{1}&\cdots&s_{N}\end{bmatrix}^{\top}$ with the element as the Ising spin variable (up, $\uparrow$, or down, $\downarrow$), $s_{i}\in\left\{\bullet=~{}\uparrow~{}\coloneqq+1,\circ=~{}\downarrow~{}\coloneqq-1\right\}$. A state $\bm{s}$ is a vertex of the $N$-dimensional hypercube, and a state transition with a single spin flip corresponds to hypercubic edge. The element of the interaction matrix $\bm{J}\in\left\{-1,0\right\}^{N\times N}$ is $J_{i,j}=J_{j,i}=-1$ when spin $i$ and $j$ antiferromagnetically interact (antialignment of spin $i$ and $j$ decreases energy) on the kagome lattice of Fig. 9(b) and $J_{i,j}=J_{j,i}=0$ otherwise. Self-interactions do not exist ($J_{i,i}=0$, $\forall i$). The external magnetic field $\bm{h}\coloneqq\begin{bmatrix}h_{1}&\cdots&h_{N}\end{bmatrix}^{\top}\in\mathbb{R}^{N}$ is assigned to be zero $\bm{0}$ in this study.
Due to the (geometrical) frustration [130] of the interaction network and associated degeneracy of ground states, the energy landscape determines the dynamics of the systems: one can obtain insights into the dynamics by visualizing the complexity of the hypercubic energy landscape. The frustration of the interaction network is quantified by the frustration function $\varPhi\left(\mathcal{C}\right)\coloneqq\operatorname{sgn}\left(\prod_{J_{i,j}\in\mathcal{C}}J_{i,j}\right)$ for a closed undirected cycle $\mathcal{C}$ in the interaction network [130]. If $\varPhi\left(\mathcal{C}\right)<0$, the cycle $\mathcal{C}$ is frustrated: any state cannot satisfy all the interactions in the cycle.
We examine the spin-ice systems through the interaction network and associated frustration. The interaction network of the one-ring system [Fig. 9(b1)] is a ring without frustration, $\varPhi\left(\mathcal{C}\right)\nless 0$, and there are two ground states: a ground state $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ and its global spin-flipped state $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\bullet&\circ\end{bmatrix}^{\top}$. The two-ring system has several ground states owing to the frustrated interaction involving spin 6, such as $\mathcal{C}=\left\{J_{6,10},J_{10,9},J_{9,6}\right\}$ and $\mathcal{C}=\left\{J_{6,3},J_{3,2},J_{2,6}\right\}$. In the three-ring system, there are more frustrated cycles containing spins 6, 9, and 10, e.g., $\mathcal{C}=\left\{J_{6,10},J_{10,9},J_{9,6}\right\}$, $\mathcal{C}=\left\{J_{6,3},J_{3,2},J_{2,6}\right\}$, $\mathcal{C}=\left\{J_{9,8},J_{8,12},J_{12,9}\right\}$, and $\mathcal{C}=\left\{J_{10,13},J_{13,11},J_{11,10}\right\}$.
VII.2 Projecting the hypercubic energy landscape of the kagome spin-ice system
For the projection of the hypercubic energy landscape, we calculate the covariance matrix [Eq. (7)] with the probability distribution of each state. The probability distribution of each state (hypercubic vertex) is determined by the canonical ensemble: $p\left(\bm{s}\right)=\frac{1}{Z}\exp\left[-\beta\mathcal{H}\left(\bm{s}\right)\right]$, where $Z\coloneqq\sum_{\bm{s}}\exp\left[-\beta\mathcal{H}\left(\bm{s}\right)\right]$ is the partition function. The temperature $T$ of the reservoir (bath) or inverse temperature $\beta$ is assigned to be $k_{\mathrm{B}}T=\frac{1}{\beta}=0.3$ in our projection. Here, $k_{\mathrm{B}}$ is the Boltzmann constant. Note that the probability distribution is symmetric with respect to the origin, $p\left(-\bm{s}\right)=p\left(\bm{s}\right)$ when the external field is zero, $\bm{h}=\bm{0}$, because the Hamiltonian has symmetry under the global spin flip, $\bm{s}\leftarrow-\bm{s}$, or $\mathbb{Z}_{2}$ symmetry, $\mathcal{H}\left(-\bm{s}\right)=-\frac{1}{2}\left(-\bm{s}\right)^{\top}\bm{J}\left(-\bm{s}\right)=-\frac{1}{2}\bm{s}^{\top}\bm{J}\bm{s}=\mathcal{H}\left(\bm{s}\right)$. This symmetry of the probability distribution makes the mean state vector the zero vector, $\bm{s}\right>=\bm{0}$ because $\bm{s}\right>=\sum_{\bm{s}}p\left(\bm{s}\right)\bm{s}=\sum_{-\bm{s}}p\left(-\bm{s}\right)\left(-\bm{s}\right)=\sum_{\bm{s}}p\left(\bm{s}\right)\left(-\bm{s}\right)=-\left<\bm{s}\right>$. Thus, the covariance matrix becomes $\bm{\varSigma}=\sum_{\bm{s}}p\left(\bm{s}\right)\bm{s}\bm{s}^{\top}$. The non-diagonal element (covariance) is $\varSigma_{i,j}=\sum_{\bm{s}}p\left(\bm{s}\right)s_{i}s_{j}$, and the diagonal element (variance) is $\varSigma_{i,i}=\sum_{\bm{s}}p\left(\bm{s}\right)s_{i}^{2}=\sum_{\bm{s}}p\left(\bm{s}\right)=1$. Therefore, the correlation between spin $i$ and $j$ is the covariance between them, $\frac{\varSigma_{i,j}}{\sqrt{\varSigma_{i,i}}\sqrt{\varSigma_{j,j}}}=\varSigma_{i,j}$. The covariance matrix becomes the correlation matrix. PCA finds the most correlated direction in the hypercubic state space.
In Figs. 10(a1), 10(a2), and 10(a3), the hypercubic energy landscapes of finite artificial kagome Ising spin systems are projected using PCA. Each state $\bm{s}$ corresponds to a vertex of the hypercube, and a state transition with a single spin flip corresponds to the edge of the hypercube. Thus, a hypercubic energy landscape is a kind of state transition diagram. For clarity of visualization, we show only the vertices with the lowest and the second-lowest energy, i.e., the ground states and the first-excited states, out of $2^{N}$ vertices (states) in Figs. 10(a1), 10(a2), and 10(a3). By visualizing hypercubic edges connecting the ground states and first-excited states, one can observe the pathways expected to be followed in the dynamics at low temperatures.
Figure 10: Hypercubic energy landscape of finite artificial spin-ice systems and interaction networks colored by the angle of the biplot basis. (a) The hypercubic energy landscape of one-ring (a1), two-ring (a2), and three-ring (a3) systems by PC1 and PC2. For clarity of the figure, only the states with the lowest and the second-lowest energy are shown. The energy of the state colors vertices (blue for the lowest energy and red for the second-lowest energy). Biplot bases are shown on the top right, and their color indicates the angle of the biplot basis. Notice that in (a2) and (a3), several biplot bases are overlapped. One example pathway from a ground state to another ground state is shown as colored hypercubic edges in (a1). The colors of these lines are the same as the color of the corresponding biplot basis. Examples of states are shown as insets in (a1), (a2), and (a3). See the caption of Fig. 9 for the definition of state visualization on the interaction network. (b) PC1 loading of one-ring (b1), two-ring (b2), and three-ring (b3) systems. The color of the bar matches the color of the corresponding biplot basis in (a). (c) Same as (b) but for PC2. (d) The Ising spin interaction networks of one-ring (d1), two-ring (d2), and three-ring (d3) systems, where the nodes are colored by the angle of the corresponding biplot basis in (a).
VII.3 The hypercubic energy landscape of the one-ring system
We first examine the one-ring system of Figs. 9(a1) and 9(b1). In Fig. 10(a1), we show the possible pathway from a ground state $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ to the globally spin-flipped state $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\bullet&\circ\end{bmatrix}^{\top}$ by connecting the vertices with the lowest energy and the second-lowest energy using hypercubic edges. (See Figs. 9(b1) or 10(d1) for labeling of spins.) In this example, the hypercubic edge corresponds to the state transition with a single spin flip. The biplot bases are shown at the top right of Fig. 10(a1), and its color indicates the angle. We emphasize one particular pathway between two ground states by coloring the edges following the biplot bases in Fig. 10(b1). We found the state transition pathway is collective: the flipping order is restricted to keep the energy low. In a one-ring system, there are $6!=720$ possible pathways from one ground state to another ground state, but the number of most probable pathways is limited, as shown in Fig. 10(a1), and it is reduced to $966\times 2^{4}\times 1=$ 96 [69]. This reduction arises from the interaction network of the system and the associated hypercubic energy landscape.
In addition, one can infer which spin is involved in which state transition by the biplot basis. One can know which spin flips in the state transition by comparing the length and angle of the hypercubic edge of interest with that of the biplot basis. The direction of the arrow of the biplot basis indicates the direction of the spin flip. In other words, if the state transition is along the direction of the arrow of the biplot basis, the spin flips from down to up. Otherwise, the spin flips from up to down. Notice that Fig. 10(a1) is the Hamming projection because the probability distribution is roughly the same as Eq. (11).
To interpret the relation between PC loading and the interaction network, we visualize the loading of PC1 and PC2 in Figs. 10(b1) and 10(c1), respectively, colored by the angle of the biplot bases in Fig. 10(a1). We confirmed that the element-wise sign of PC1 corresponds to the ground state [Fig. 10(b1)]. The element-wise sign of PC2 loading corresponds to the state perpendicular to the ground state, which has a Hamming distance of $\frac{N}{2}$ from the ground state [Fig. 10(c1)].
The interaction network in Fig. 10(d1), where nodes are colored by the angle of the biplot basis, shows the correlation captured by the first two PC loadings. Spins 1 and 5 have the same color, and spin 3, interacting with both spins 1 and 5, has the color complementary to that of spins 1 and 5. We can argue the same for spins 2, 4, and 6. Hence, the resulting hypercubic energy landscape is drawn to emphasize the correlation arising from the interaction.
VII.4 The hypercubic energy landscape of the two-ring system
The hypercubic energy landscape of the two-ring system, which has several ground states, is illustrated in Fig. 10(a2). We show six ground states and pathways between them in a two-ring system with a biplot basis on the top right. Similar to the one-ring system in Fig. 10(a1), the most probable pathways between the ground states are shown by connecting the vertices with the lowest and the second-lowest energy. We discuss two pathways connecting the ground states as examples.
The first example involves two ground states on the left side (around $\begin{bmatrix}-3&0\end{bmatrix}^{\top}$), i.e., $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\bullet&\circ&\circ&\bullet&\circ&\bullet\end{bmatrix}^{\top}$ and $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\circ&\circ&\circ&\bullet&\circ&\bullet\end{bmatrix}^{\top}$ which are connected by the hypercubic edge involving the flip of spin 6, indicating that spin 6 is expected to fluctuate on the left side of the landscape. One of the interactions in pairs of $J_{6,3}$ and $J_{6,10}$, or $J_{6,2}$ and $J_{6,9}$, is unsatisfied with any state of spin 6. This frustration of the interaction network cancels the local field to spin 6, i.e., $\sum_{j=1}^{11}J_{6,j}s_{j}=0$. This is the reason why spin 6 does not have an energetically favored state.
The second example is the transition pathway from the ground state on the upper left side (around $\begin{bmatrix}-3&0.3\end{bmatrix}^{\top}$), to the ground state on the top (around $\begin{bmatrix}0&3.2\end{bmatrix}^{\top}$), i.e., state transition from $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\bullet&\circ&\circ&\bullet&\circ&\bullet\end{bmatrix}^{\top}$ to $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$, which requires the flipping of spins 1, 2, 5, 8, and 9 [see biplot bases of Fig. 10(a2) and color of Figs. 10(b2), 10(c2), and 10(d2)]. This projection of the hypercube indicates that state transition happens with the correlated spin flips on the left-half part of the system (spins 1, 2, 5, 8, and 9). Among the left-half part of the system, spins 2 and 9 are more likely to flip at the beginning of the pathway. This difference arises from the unsatisfied interactions: $J_{3,9}$ and $J_{6,2}$. Spins 2 and 9 receive weaker local fields than spins 1, 5, and 8 because they have interactions involving spins 3, 6, and 10. Thus, spins 2 and 9 are expected to flip more than spins 1, 5, and 8. Because of the frustrated interactions involving spins 3, 6, and 10, spins 2 or 9 are the most probable to flip first. Then, the flips happen in the order of spins 2, 1, 5, 8, and 9 if spin 2 flips first, or in the order of spins 9, 8, 5, 1, and 2 if spin 9 flips first. Similar to the one-ring system in Fig. 10(a1), the number of most probable pathways is limited, but the constraints are stricter and fewer pathways are possible.
To deepen our understanding of PC loading and interaction networks, we show the loading of PC1 and PC2 in Figs. 10(b2) and 10(c2), coloring them by the angle of the biplot bases in Fig. 10(a2). Again, we confirmed that the sign of PC1 loading [Fig. 10(b2)] corresponds to the ground state with the largest magnitude of the PC1 score, but spin 6 is not determined. Thus, the frustrated dynamics of spin 6 are captured by PC1. The element-wise sign of PC2 loading [Fig. 10(c2)] corresponds to the ground state with the largest magnitude of the PC2 score. Notice that the sixth element of PC2 loading has a slightly higher magnitude than the others, indicating the importance of spin 6. If spin 6 flips when the system is in the ground state $\operatorname{sgn}\left(\bm{u}_{2}\right)$, the energy of the system increases more than when another spin flips.
The interaction network with nodes colored by the angle of the biplot bases in Fig. 10(d2) shows the inter-spin correlation captured by the biplot bases. The biplot bases capture the (anti)correlation of spins 1, 2, 5, 8, and 9 [yellow and blue arrows in Fig. 10(a2) top right] which corresponds to the left ring of the interaction network in Fig. 10(d2). The biplot bases also capturs the (anti)correlation of spins 3, 4, 7, 10, and 11 [green and purple arrows in Fig. 10(a2) top right] which corresponds to the right ring of the interaction network in Fig. 10(d2). The hypercubic energy landscape using PCA provides insight into the dynamics through the correlation of spins.
VII.5 The hypercubic energy landscape of the three-ring system
More ground states and probable pathways emerge in the hypercubic energy landscape of a three-ring system [Fig. 10(a3)]. There are eight groups of connected ground states; two of them around $\begin{bmatrix}-3&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}3&0\end{bmatrix}^{\top}$ consist of six ground states, and the rest consist of two ground states.
Unlike the one-ring and two-ring systems, the structure of pathways is more complex, but we can still infer the state transition from the projection. For example, the pathway from the ground state $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\circ&\circ&\circ&\circ&\bullet&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ (around $\begin{bmatrix}-3&0.9\end{bmatrix}^{\top}$) to $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ (around $\begin{bmatrix}-1.5&3.2\end{bmatrix}^{\top}$) is a transition with the correlated spin flips of spins located in the lower right hexagon (spins 3, 4, 7, and 11). Pathways parallel to this pathway involve the same spin flips. From the biplot bases on the top right of Fig. 10(a3), one can generally know that by following the angle of each biplot basis, the transitions involve spins on the top hexagon (spins 12, 13, 14, and 15) being flipped: the pathways are lines from the left top to the right bottom, which are parallel to the biplot bases of spins 12, 13, 14, and 15. Likewise, horizontal pathways are used when spins in the lower left hexagon (spins 1, 2, 5, and 8) are flipped, and vertical pathways are followed when spins in the inverted triangle at the center (spins 6, 9, and 10) are flipped. Similar to the one-ring and two-ring systems, the pathway is a collective flip of spins reflecting the interaction network. Notice that this projection of the hypercubic energy landscape [Fig. 10(a3)] is a partial Hamming projection: the PC1 score is equivalent to the Hamming distance from the ground state on the left side (around $\begin{bmatrix}-3&0\end{bmatrix}^{\top}$) or the right side (around $\begin{bmatrix}3&0\end{bmatrix}^{\top}$) but without considering spins 6, 9, and 10.
To understand the relation between PC loading and the interaction network, in particular, the relation between PC loading and the frustration of the system, we color the loading of PC1 and PC2 in Figs. 10(b3) and 10(c3), following the angle of the biplot bases in Fig. 10(a3). We confirmed that the element-wise sign of PC1 [Fig. 10(b3)] corresponds to the ground state group containing six ground states, but the states of spins 6, 9, and 10 are not determined. These spins involve a frustrated cycle of the interaction network. PC2 loading in Fig. 10(c3) shows that spins 1, 2, 5, and 8 do not contribute, and those spins belong to the lower left ring. Moreover, PC2 loading of spins 6, 9, and 10 has slightly higher or lower magnitudes than the others, indicating the uncommon contribution of spins 6, 9, and 10, suggesting the frustration-related correlation.
These insights are consistent with the interaction network with colored nodes in Fig. 10(d3): the PC loading captures the correlation arising from the interaction network. Nevertheless, spins 6 and 9 have exactly the same color: their biplot bases are the same, but it can be another combination of spins such as 6 and 10. This breaks the symmetry of the system—PC2 ignores the lower left hexagon of the interaction network even though the system has the same other structure—and we examine how the PC loading reflects the symmetry of the system.
Figure 11: Fraction of explained variance by each PC of the hypercubic energy landscapes in Fig. 10. (a1) One-ring system. (a2) Two-ring system. (a3) Three-ring system.
VII.6 Fraction of explained variance
We investigate the fraction of explained variance by the PCs of the hypercubic energy landscape in Fig. 11. to see how the frustration of the system influences the PCs. For the one-ring system [Fig. 11(a1)], PC1 dominates the explained variance by the PCs, which is consistent with PC1 being proportional to the ground state. Turning to the two-ring and three-ring systems [Figs. 11(a2) and 11(a3)], the first PCs explain less than 90% of the variance. This decrease in the explained variance by the first component arises from the degeneracy of the ground states. PC2 explains the variance of degenerate ground states. For the three-ring system [Fig. 11(a3)], the first three PCs are required to explain more than 80% of the variance, and the second and third fractions of explained variance are the same, indicating that PC2 and PC3 share the regularity: the symmetry of the system.
Figure 12: Same as Figs. 10(a3), 10(b3), 10(c3), and 10(d3) but by different PCs. (a) The hypercubic energy landscape of the three-ring system by PC1 and PC3. (b) PC1 loading of (a) and the color of the bar matches the color of the corresponding biplot basis in (a). (c) Same as (b) but for PC3. (d) The Ising spin interaction networks of the three-ring systems, where the nodes are colored by the angle of the corresponding biplot basis in (a). (e–h) Same as (a–d) but by PC2 and PC3. Notice that in (a) and (e), several biplot bases are overlapped.
VII.7 The hypercubic energy landscape of the three-ring system through PC3
The fraction of explained variance of the three-ring system [Fig. 11(a3)] leads us to visualize the hypercubic energy landscapes of three-ring101010 See the Supplemental Material [105] for hypercubic energy landscapes of two-ring systems using PC3. systems using PC3. In Fig. 12(a), the hypercubic energy landscape of three-ring systems, the same as Fig. 10(a3), is shown but by PC1 and PC3. We find that with PC1 and PC3, unlike the projection by PC1 and PC2, the shape of the hypercubic energy landscape looks similar to a parallelogram, not a hexagon. The biplot bases on the top right of Fig. 12(a) are only six arrows despite the system having 15 spins, indicating numerous overlaps of the biplot bases. To check which spins are overlapped, we show the loading of PC1 and PC3 in Figs. 12(b) and 12(c), and the interaction network in Fig. 12(d), coloring them by the angle of the biplot basis. PC3 loading [Fig. 12(c)], in particular, has a nonuniform magnitude of the elements; the magnitude of the elements of spins 1, 2, 5, 6, 8, and 9 are larger than the others. Those spins belong to the lower left hexagon of the interaction network in Fig. 12(d). The interaction network colored by the angle of the biplot basis in Fig. 12(d) illustrates how the biplot bases captures the correlation of spins. As indicated by PC3 loading [Fig. 12(c)], the spins on the lower left hexagon (spins 1, 2, 5, 6, 8, and 9) form the (anti)correlated group. The rest of the two rings form a large (anti)correlated group. PC3 loading creates the hypercubic energy landscape, emphasizing the lower left hexagon of the interaction network.
We further investigate the hypercubic energy landscape of the three-ring system by PC2 and PC3 to understand the symmetry of the system. The hypercubic energy landscape of the three-ring system by PC2 and PC3 [Fig. 12(e)], looks like a hexagon, similar to the projection by PC1 and PC2. Nevertheless, the landscape is more symmetric and regular than that by PC1 and PC2. Around the origin, the connected six ground states also form a regular hexagon, but they are overlapping with the other group of six ground states, cf. Fig. 10(a3). The angles of the biplot bases are almost uniform (with some overlapping), and the lengths of them are nearly the same, resulting in a symmetric and regular hypercubic energy landscape; this projection is the same kind as the isometric projection, but that of a nine-dimensional hypercube.
The loading of PC2 and PC3 is shown in Figs. 12(f) and 12(g), with the node color indicating the angle of the biplot basis. As we mentioned in Sec. VII.5, the PC2 loading emphasizes contributions from spins 6, 9, and 10 but ignores spins 1, 2, 5, and 8. On the contrary, PC3 loading has a higher magnitude for spins 1, 2, 5, and 8 in addition to spins 6 and 9, but ignores spin 10. Spins belonging to other hexagons contribute equally to the PC2 and PC3 loading. Therefore, as shown in the biplot basis of Fig. 12(a), the contributions by PC2 and PC3 are complementary, and the resulting hypercubic energy landscape shows the symmetry of the state space reflecting the interaction network.
The interaction network colored by the angle of the biplot basis in Fig. 12(h) supports our interpretation: the coloring is symmetric on the interaction network. For example, there is a stripe pattern along the outer interaction cycle (spins 1, 2, 3, 4, 7, 11, 13, 15, 14, 12, 8, and 5). The smaller interaction cycle consisting of six spins (such as spins 1, 2, 6, 9, 8, and 5) also exhibits the stripe pattern. There is a red-green-blue stripe pattern in the interaction cycle with three spins (such as spins 6, 10, and 9). The stripe pattern in all sizes of interaction cycles indicates that the three-ring system has a collective mode involving the whole system, and those are hierarchical: the stripe pattern of the outer cycle arises from the stripe pattern of the smaller cycle. This non-local correlation is captured by the PC2 and PC3 loading as the symmetry of the interaction network. With an appropriate combination of PC loading, such as PCs sharing the same fraction of explained variance, PCA can capture the symmetry of the system.
VIII Conclusions
The hypercubic representation of binary state space is a powerful tool to reveal the high-dimensional structures of binary data in various sciences, yet as R. P. Feynman once pointed out [131], “unfortunately our brains can’t visualize” the high-dimensional hypercube. Conventional methods of projection have both advantages and disadvantages, and there is a demand for informative and practical methods. In this study, we suggest several orthogonal projection methods to obtain reproducible, interpretable, and automatic visualizations of hypercubes, concluding that PCA is a suitable method for this purpose. We merge ideas from high-dimensional geometry and statistics (unsupervised machine learning), and apply the projections of hypercubes to understand statistical mechanical models. Extending our work to other linear dimensionality reduction methods is straightforward if one obtains a biplot.
Our contributions are summarized as follows: (1) We introduce the Hamming and fractal projection of hypercubes in Sec. II. (2) We propose interpreting the biplot bases of PCA as the bases of the orthogonal projection of hypercubes in Sec. III. (3) We find that the PC1 loading of PCA on hypercubes is equivalent to the weighted hypercubic vertices in Sec. IV, and validate it analytically and numerically in Sec. V. (4) We reveal numerically and analytically that the quality of the orthogonal projection of hypercubes tends to be worse around the origin of the projected space in Sec. VI. (5) We visualize the hypercubic energy landscapes of finite artificial spin-ice systems and extract physical interpretations of them in Sec. VII.
Although we demonstrate that PCA produces reproducible, interpretable, and automatic projections of hypercubes, the remaining challenge is the overlapping of the hypercubic vertices in the projection. As we show numerically and analytically in Sec. VI, PCA locates numerous hypercubic vertices around the origin of the projected space, and those overlapped hypercubic vertices can increase the error of the projection. We, however, analytically know these limitations of PCA beforehand: we can interpret the resulting plot while grasping the limitations, unlike other nonlinear dimensionality reduction methods, where such analytical properties are not obvious and might be more difficult to understand. When we create the hypercubic energy landscape in Sec. VII, we avoid the overlapping problem by projecting a limited number of hypercubic vertices. With a similar strategy, one might bypass this difficulty in other applications.
H. S. M. Coxeter mentioned that illustrations of high-dimensional objects have psychological and artistic merit [9]. In this study, we initiate unveiling the new merit—or possibility—of projecting high-dimensional hypercubes: interpreting binary data through data-driven visualization. As Anscombe’s quartet [132] shows hidden patterns not in statistics but in graphs, informative orthogonal projection using PCA might lead to fresh understanding, interpretation, and discovery [133, 134] in high-dimensional binary data across sciences.
Calculations and visualizations of this work were performed using open-source python [135] libraries: matplotlib [136], networkx [137], numpy [138], and scipy [139]. All codes will be available online [140].
Acknowledgements.
Special thanks go to Takashi Horiike, Kennichi Hagiwara, Nozomu Imai, Takumi Nishina, and Hiroto Murata for fruitful discussions at the early stage of this research. We thank valuable comments on this manuscript from Mizuho Aoki, Johanna von Drachenfels, Anton G. Golles, and Henri Schmidt. This work was supported by KAKENHI Grants 22H00406 and 24H00061 of the Japan Society for the Promotion of Science.
Y.H. performed conceptualization, formal analysis, software development, validation, investigation, data curation, and visualization; Y.H. and S.F. developed the methodology; S.F. provided computational resources; Y.H. wrote the original draft; and Y.H. and S.F. reviewed and edited the manuscript.
Appendix A Constructing hypercubes
A hypercube is constructed as follows [9, 141, 12]. Suppose one has a zero-dimensional point, one obtains a one-dimensional line by moving the point in a direction. By moving the line in a direction but not along its line, one can obtain a two-dimensional parallelogram. Again, moving the parallelogram in a direction but not along its plane, one creates a three-dimensional parallelepiped. By repeating this translation process $N$ times, one obtains an $N$-dimensional parallelepiped or a parallelotope with $2^{N}$ vertices and $\frac{2^{N}N}{2}$ edges.111111 Each of the $2^{N}$ vertices has $N$ edges. To avoid double counting, one needs to divide them by two. The unit vectors of each dimension correspond to the direction of the translation, i.e., edges created by the translation are parallel to the directions of unit vectors.
Some special objects of parallelotopes are known [9]. An orthotope is a special parallelotope where the unit vectors of each dimension are mutually perpendicular. A hypercube (also written hyper-cube) or measure polytope is a type of orthotope where all unit vectors of each dimension share the same length. An $N$-dimensional hypercube is labeled as $\gamma_{N}$.
Appendix B Formulations of PCA
In this Appendix, we briefly review four formulations of PCA [102, 103, 121, 122, 123]. Our review includes the well-known two formulations: the maximum projection variance formulation [101] and the minimum reconstruction error formulation [100]. Then, the equivalence of classical MDS and PCA is reintroduced [120], and we present the minimum inner-product formulation of PCA [123].
Throughout this Appendix B, we use an $N$-dimensional vector with real elements $\bm{x}\in\mathbb{R}^{N}$ as the data point in the original space. In our main text, we use the Ising state vector $\bm{s}\in\left\{+1,-1\right\}^{N}$ as hypercubic vertices (states or data points), so replace $\bm{x}$ with $\bm{s}$ if readers consider specifically the projections of hypercubic vertices.
B.1 Maximum projection variance formulation
[…]
$L\coloneqq$ $\sum_{\bm{x}}p\left(\bm{x}\right)\left[r_{1}\left(\bm{x}\right)-\left<r_{1}\left(\bm{x}\right)\right>\right]^{2}$ (47)
$=$ $\sum_{\bm{x}}p\left(\bm{x}\right)\left[\left(\bm{x}-\left<\bm{x}\right>\right)^{\top}\bm{u}_{1}\right]^{2}$
[…]
$=$ $\bm{u}_{1}^{\top}\bm{\varSigma}\bm{u}_{1},$ (48)
[…]
$\mathcal{L}\coloneqq\bm{u}_{1}^{\top}\bm{\varSigma}\bm{u}_{1}+\lambda_{1}\left(1-\bm{u}_{1}^{\top}\bm{u}_{1}\right),$ (49)
where $\lambda_{1}$ is the Lagrange multiplier. Deriving $\mathcal{L}$ with respect to $\bm{u}_{1}$ and setting it to be the zero vector,
$\bm{0}=\frac{\partial}{\partial\bm{u}_{1}}\mathcal{L}=2\bm{\varSigma}\bm{u}_{1}-2\lambda_{1}\bm{u}_{1},$ (50)
one obtains $\bm{\varSigma}\bm{u}_{1}=\lambda_{1}\bm{u}_{1}$ and the variance of the projected data set is $\bm{u}_{1}^{\top}\bm{\varSigma}\bm{u}_{1}=\lambda_{1}.$ The eigenvector $\bm{u}_{1}$ is the PC1 explaining the variance $\lambda_{1}$ which is maximized.
The further PCs are obtained by an incremental procedure while keeping the orthogonality to all the previous PCs. For example, PC2 is obtained by the same procedure but with the normalization constraint $\bm{u}_{2}^{\top}\bm{u}_{2}=1$ and the orthogonality constraint to PC1 $\bm{u}_{1}^{\top}\bm{u}_{2}=0$. The Lagrange function is
$\mathcal{L}\coloneqq\bm{u}_{2}^{\top}\bm{\varSigma}\bm{u}_{2}+\lambda_{2}\left(1-\bm{u}_{2}^{\top}\bm{u}_{2}\right)+\eta\bm{u}_{1}^{\top}\bm{u}_{2},$ (51)
where $\lambda_{2}$ and $\eta$ are Lagrange multipliers. Similarly to PC1, deriving $\mathcal{L}$ with respect to $\bm{u}_{2}$ and setting it to be the zero vector, one obtains
$\bm{0}=\frac{\partial}{\partial\bm{u}_{2}}\mathcal{L}=2\bm{\varSigma}\bm{u}_{2}-2\lambda_{2}\bm{u}_{2}+\eta\bm{u}_{1}.$ (52)
Multiplying $\bm{u}_{1}^{\top}$ to both sides of the equation from the left side yields121212 $\bm{u}_{1}^{\top}\bm{\varSigma}\bm{u}_{2}=\bm{u}_{2}^{\top}\bm{\varSigma}\bm{u}_{1}=\bm{u}_{2}^{\top}\lambda_{1}\bm{u}_{1}=\lambda_{1}\bm{u}_{2}^{\top}\bm{u}_{1}=0$ because of the orthogonality between $\bm{u}_{1}$ and $\bm{u}_{2}$, $\bm{u}_{1}^{\top}\bm{u}_{2}=0$ [102]. $0=\bm{u}_{1}^{\top}\bm{\varSigma}\bm{u}_{2}-\lambda_{2}\bm{u}_{1}^{\top}\bm{u}_{2}+\eta\bm{u}_{1}^{\top}\bm{u}_{1}=\eta\bm{u}_{1}^{\top}\bm{u}_{1}=\eta$ resulting in $\eta=0$. Thus, we obtain $\bm{\varSigma}\bm{u}_{2}=\lambda_{2}\bm{u}_{2}$ and $\bm{u}_{2}$ is also the eigenvector of the covariance matrix $\bm{\varSigma}$, explaining variance $\bm{u}_{2}^{\top}\bm{\varSigma}\bm{u}_{2}=\lambda_{2}.$ The same procedure is applied to obtain further eigenvectors (PCs).
B.2 Minimum reconstruction error formulation
[…]
$L\coloneqq$ $\sum_{\bm{x}}p\left(\bm{x}\right)\left|\bm{x}-\tilde{\bm{x}}\right|^{2}$ (53)
$=$ $\sum_{\bm{x}}p\left(\bm{x}\right)\left|\sum_{i=1}^{n}\left[r_{i}\left(\bm{x}\right)-f_{i}\left(\bm{x}\right)\right]\bm{u}_{i}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\left.+\sum_{i=n+1}^{N}\left[r_{i}\left(\bm{x}\right)-g_{i}\right]\bm{u}_{i}\right|^{2}$
$=$ $\sum_{\bm{x}}p\left(\bm{x}\right)\sum_{i=1}^{n}\left[r_{i}\left(\bm{x}\right)-f_{i}\left(\bm{x}\right)\right]^{2}$
$+\sum_{\bm{x}}p\left(\bm{x}\right)\sum_{i=n+1}^{N}\left[r_{i}\left(\bm{x}\right)-g_{i}\right]^{2},$ (54)
which is minimized by choosing the appropriate $\left\{f_{i}\left(\bm{x}\right)\right\}_{i=1}^{n}$, $\left\{g_{i}\right\}_{i=n+1}^{N}$, and $\left\{\bm{u}_{i}\right\}_{i=1}^{N}$.
We first derive $L$ with respect to $f_{i}\left(\bm{x}\right)$, resulting in
$0=\frac{\partial}{\partial f_{i}\left(\bm{x}\right)}L=-2\sum_{\bm{x}}p\left(\bm{x}\right)\left[r_{i}\left(\bm{x}\right)-f_{i}\left(\bm{x}\right)\right].$ (55)
We obtain $\sum_{\bm{x}}p\left(\bm{x}\right)f_{i}\left(\bm{x}\right)=\sum_{\bm{x}}p\left(\bm{x}\right)r_{i}\left(\bm{x}\right)$ which is equivalent to $f_{i}\left(\bm{x}\right)=r_{i}\left(\bm{x}\right)$. Then, we derive $L$ with respect to $g_{i}$, getting
$0=\frac{\partial}{\partial g_{i}}L=-2\sum_{\bm{x}}p\left(\bm{x}\right)\left[r_{i}\left(\bm{x}\right)-g_{i}\right].$ (56)
The solution is $\sum_{\bm{x}}p\left(\bm{x}\right)g_{i}=\sum_{\bm{x}}p\left(\bm{x}\right)r_{i}\left(\bm{x}\right)$ which can further be simplified as $g_{i}=\left<r_{i}\left(\bm{x}\right)\right>.$ Then, the reconstruction error becomes
$L=$ $\sum_{\bm{x}}p\left(\bm{x}\right)\sum_{i=n+1}^{N}\left[r_{i}\left(\bm{x}\right)-\left<r_{i}\left(\bm{x}\right)\right>\right]^{2}$
$=$ $\sum_{\bm{x}}p\left(\bm{x}\right)\sum_{i=n+1}^{N}\left[\left(\bm{x}-\left<\bm{x}\right>\right)^{\top}\bm{u}_{i}\right]^{2}$
[…]
$=$ $\sum_{i=n+1}^{N}\bm{u}_{i}^{\top}\bm{\varSigma}\bm{u}_{i}.$ (57)
We then minimize $L$ by choosing appropriate $\left\{\bm{u}_{i}\right\}_{i=n+1}^{N}$ by the method of Lagrange multipliers. The Lagrange function is
$\mathcal{L}\coloneqq\underbrace{\sum_{i=n+1}^{N}\bm{u}_{i}^{\top}\bm{\varSigma}\bm{u}_{i}}_{L}+\sum_{i=n+1}^{N}\lambda_{i}\left(1-\bm{u}_{i}^{\top}\bm{u}_{i}\right).$ (58)
[…]
B.3 Distance, similarity, or overlap preserving formulation
In this subsection, we briefly review the pairwise distance-preserving formulation of PCA via classical MDS [120, 121, 123]. The goal of classical MDS is to find a set of points in a lower-dimensional space that preserves the original pairwise distances (or similarity). To achieve this, we rewrite the squared Euclidean distance using the inner product. Then, by performing eigendecomposition of the inner product matrix, we obtain the eigenvectors, which are the desired lower-dimensional coordinates. Finally, we show that this classical MDS is equivalent to PCA. We first consider the classical MDS with the squared Euclidean distance then try to formulate it with the probability of data points.
B.3.1 Classical MDS
A popular distance measure is the squared Euclidean distance. Suppose we have $M$ data points $\left\{\bm{x}_{i}\right\}_{i=1}^{M}$. We assume $M>N$, i.e., the number of data points is larger than the dimension of the data. We consider the Euclidean distance matrix $\bm{D}\in\mathbb{R}_{\geq 0}^{M\times M}$ but with the squared elements $D_{i,j}\coloneqq D_{\mathrm{E}}^{2}\left(\bm{x}_{i},\bm{x}_{j}\right)$ between the $i$th and $j$th data points. The squared Euclidean distance is
$D_{i,j}=$ $\left|\bm{x}_{i}-\bm{x}_{j}\right|^{2}$
$=$ $\left(\bm{x}_{i}-\bm{x}_{j}\right)^{\top}\left(\bm{x}_{i}-\bm{x}_{j}\right)$
$=$ $\bm{x}_{i}^{\top}\bm{x}_{i}-2\bm{x}_{i}^{\top}\bm{x}_{j}+\bm{x}_{j}^{\top}\bm{x}_{j}$
$=$ $G_{i,i}-2G_{i,j}+G_{j,j},$ (59)
where we introduce the similarity (inner product) matrix or the Gram matrix $\bm{G}\in\mathbb{R}^{M\times M}$ which has the inner product as its element $G_{i,j}\coloneqq\bm{x}_{i}^{\top}\bm{x}_{j}$. Notice that the element of the Gram matrix is proportional to the overlap, i.e., $G_{i,j}=NQ\left(\bm{s}_{i},\bm{s}_{j}\right)$ if the data points are Ising spin states $\bm{x}\rightarrow\bm{s}$. We can write the element as $D_{i,j}=g_{i}-2G_{i,j}+g_{j}.$ Here, $g_{i}\coloneqq G_{i,i}$ is the $i$th diagonal element of the Gram matrix $\bm{G}$ and $\bm{g}=\operatorname{diag}\left(\bm{G}\right)$ is the vector of diagonal elements of the Gram matrix. Then, Eq. (59) becomes, in matrix form,
$\bm{D}=\bm{g}\bm{1}^{\top}-2\bm{G}+\bm{1}\bm{g}^{\top},$ (60)
where $\bm{1}\coloneqq\begin{bmatrix}1&\cdots&1\end{bmatrix}^{\top}\in\left\{1\right\}^{M}$ is the vector of $M$ ones. We then introduce the centering matrix as $\bm{C}\coloneqq\bm{I}-\frac{1}{M}\bm{1}\bm{1}^{\top},$ where $\bm{I}=\operatorname{diag}\left(\bm{1}\right)$ is the $M\times M$ identity matrix. Notice that the centering matrix is symmetric $\bm{C}^{\top}=\bm{C}$ and idempotent $\bm{C}^{2}=\bm{I}-\frac{2}{M}\bm{1}\bm{1}^{\top}+\frac{1}{M}\bm{1}\bm{1}^{\top}=\bm{C}$. The centering matrix subtracts the mean of the column (row) of a matrix when it is multiplied from the left (right) side of the matrix. Next, we reduce the mean of both row and column of the squared Euclidean distance matrix, i.e., we double-center the squared Euclidean distance matrix, resulting in
$\bm{C}\bm{D}\bm{C}=$ $\bm{C}\left(\bm{g}\bm{1}^{\top}-2\bm{G}+\bm{1}\bm{g}^{\top}\right)\bm{C}$
$=$ $-2\bm{C}\bm{G}\bm{C},$ (61)
because $\bm{1}^{\top}\bm{C}=\bm{0}^{\top}$ and $\bm{C}\bm{1}=\bm{0}$. Here, $\bm{0}\coloneqq\begin{bmatrix}0&\cdots&0\end{bmatrix}^{\top}\in\left\{0\right\}^{M}$ is the vector of $M$ zeros. Thus, when we double-center the Euclidean distance matrix, we obtain the double-centered Gram matrix, i.e., the distance matrix becomes the similarity matrix.
We perform eigendecomposition of the Gram matrix
$\bm{G}=\bm{W}\bm{\varOmega}\bm{W}^{\top},$ (62)
where $\bm{W}\coloneqq\begin{bmatrix}\bm{w}_{1}&\cdots&\bm{w}_{M}\end{bmatrix}\in\mathbb{R}^{M\times M}$ is the matrix with columns being eigenvectors $\{\bm{w}_{i}\}$ and $\bm{\varOmega}\coloneqq\operatorname{diag}\left(\omega_{1},\ldots,\omega_{M}\right)\in\mathbb{R}_{\geq 0}^{M\times M}$ is the diagonal matrix of eigenvalues $\left\{\omega_{i}\right\}$ in descending order $\omega_{1}\geq\cdots\geq\omega_{M}\geq 0$. Noticing that the Gram matrix is the inner product matrix, we can express the Gram matrix as a multiplication of matrices
$\bm{G}\coloneqq\bm{X}^{\top}\bm{X},$ (63)
where $\bm{X}\coloneqq\begin{bmatrix}\bm{x}_{1}&\cdots&\bm{x}_{M}\end{bmatrix}\in\mathbb{R}^{N\times M}$ is the data matrix, with each column being a data point. Then, the rank of the Gram matrix becomes $\operatorname{rank}\left(\bm{G}\right)=\operatorname{rank}\left(\bm{X}^{\top}\bm{X}\right)=\operatorname{rank}\left(\bm{X}\right)=N$. Because the Gram matrix $\bm{G}$ is symmetric and positive semi-definite of rank $N$, it has $N$ non-negative eigenvalues and $M-N$ zero eigenvalues: $\omega_{1}\geq\cdots\geq\omega_{N}>\omega_{N+1}=\cdots=\omega_{M}=0$. We then rewrite the eigendecomposition of the Gram matrix [Eq. (62)] as
$\bm{G}=$ $\tilde{\bm{W}}\tilde{\bm{\varOmega}}\tilde{\bm{W}}^{\top}$ (64)
$=$ $\tilde{\bm{W}}\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{W}}^{\top},$ (65)
with reduced eigenvalue diagonal matrix $\tilde{\bm{\varOmega}}\coloneqq\operatorname{diag}\left(\omega_{1},\ldots,\omega_{N}\right)\in\mathbb{R}_{\geq 0}^{N\times N}$ and reduced eigenvector matrix $\tilde{\bm{W}}\coloneqq\begin{bmatrix}\bm{w}_{1}&\cdots&\bm{w}_{N}\end{bmatrix}\in\mathbb{R}^{M\times N}$ which exclude the zero eigenvalues and corresponding eigenvectors, respectively.
Calculating the Gram matrix with the centered data matrix, $\bm{X}\leftarrow\bm{X}\bm{C}$, corresponds to the double centering of the distance matrix
$\bm{G}=$ $\bm{C}\bm{X}^{\top}\bm{X}\bm{C}$ (66)
$=$ $-\frac{1}{2}\bm{C}\bm{D}\bm{C}.$
By comparing Eqs. (65) and (66), one sees that the centered data matrix is expressed as
$\bm{X}\bm{C}=\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{W}}^{\top}.$ (67)
To approximate the original data points by the lower-dimensional coordinates while maximally preserving the pairwise distances, one can use the desired number of the first rows of $\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{W}}^{\top}$. For example, if one wants to approximate the original $N$-dimensional data point by two-dimensional coordinates, classical MDS provides a set of scaled coordinates $\bm{R}=\begin{bmatrix}\bm{r}_{1}&\cdots&\bm{r}_{M}\end{bmatrix}=\begin{bmatrix}\sqrt{\omega_{1}}\bm{w}_{1}&\sqrt{\omega_{2}}\bm{w}_{2}\end{bmatrix}^{\top}\in\mathbb{R}^{2\times M},$ where $\bm{r}_{i}$ is the two-dimensional scaled coordinate of the $i$th data point. Replacing original coordinates with principal coordinates is also called principal coordinate analysis (PCO or PCoA) [120].
Numerically, the classical MDS is performed by four steps: (1) calculating the squared Euclidean distance matrix, (2) performing double centering of the distance matrix, (3) eigendecomposition of the double-centered distance matrix, and (4) obtaining the scaled (principal) coordinates from the eigenvalues and eigenvectors.
B.3.2 Classical MDS is equivalent to PCA
Here, we show that obtaining the scaled coordinates by classical MDS is equivalent to PCA. From Eq. (64), the eigenvalues and eigenvectors of the double-centered Gram matrix satisfy
$\underbrace{\bm{C}\bm{X}^{\top}\bm{X}\bm{C}}_{\bm{G}}\bm{w}_{i}=\omega_{i}\bm{w}_{i},$ (68)
[…]
$\underbrace{\frac{1}{M}\bm{X}\bm{C}\bm{X}^{\top}}_{\bm{\varSigma}}\left(\bm{X}\bm{C}\bm{w}_{i}\right)=\frac{\omega_{i}}{M}\left(\bm{X}\bm{C}\bm{w}_{i}\right).$ (69)
Notice $\bm{C}^{\top}=\bm{C}$ and $\bm{C}^{2}=\bm{C}$. We introduce the covariance matrix in the matrix form
$\bm{\varSigma}\coloneqq\frac{1}{M}\left(\bm{X}\bm{C}\right)\left(\bm{X}\bm{C}\right)^{\top}=\frac{1}{M}\bm{X}\bm{C}\bm{X}^{\top}.$ (70)
Note that PCA performs eigendecomposition of the covariance matrix $\bm{\varSigma}\bm{u}_{i}=\lambda_{i}\bm{u}_{i}$. Thus, from Eq. (69), the eigenvalues of the covariance matrix are proportional to those of the Gram matrix,
$\lambda_{i}=\frac{\omega_{i}}{M},$ (71)
and the eigenvector of the covariance matrix is expressed with the eigenvector of the Gram matrix as
$\bm{u}_{i}=\frac{1}{\sqrt{\omega_{i}}}\bm{X}\bm{C}\bm{w}_{i}.$ (72)
Following Eq. (68), or $\bm{w}_{i}^{\top}\bm{C}\bm{X}^{\top}\bm{X}\bm{C}\bm{w}_{i}=\omega_{i}$, the normalization factor $\frac{1}{\sqrt{\omega_{i}}}$ is introduced. The PC score by PC loading $\bm{u}_{i}$ and the $i$th element of principal coordinates $\sqrt{\omega_{i}}\bm{w}_{i}^{\top}$ are equivalent:
$\bm{u}_{i}^{\top}\bm{X}\bm{C}=$ $\frac{1}{\sqrt{\omega_{i}}}\bm{w}_{i}^{\top}\bm{C}\bm{X}^{\top}\bm{X}\bm{C}$
$=$ $\sqrt{\omega_{i}}\bm{w}_{i}^{\top}.$ (73)
Thus, the results of classical MDS are equivalent to those of PCA. This derivation of PCA provides us with a new understanding of PCA through pairwise distance or inner product.
B.3.3 Weighted classical MDS as PCA
We then extend classical MDS with the probability of data points. Our goal is to find the covariance matrix,
$\bm{\varSigma}\coloneqq\bm{X}\acute{\bm{C}}\bm{P}\acute{\bm{C}}^{\top}\bm{X}^{\top}=\left(\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\right)\left(\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\right)^{\top},$ (74)
from the Gram matrix. Here, $\bm{P}\coloneqq\operatorname{diag}\left(\bm{p}\right)\in\mathbb{R}_{\geq 0}^{M\times M}$ is the probability matrix and $\bm{p}\in\left[0,1\right]^{M}$ is the vector of probabilities, satisfying $\operatorname{tr}\left(\bm{P}\right)=1$. Because of the probabilities, the centering matrix is modified as $\acute{\bm{C}}\coloneqq\bm{I}-\bm{1}\bm{p}^{\top}=\bm{I}-\bm{1}\bm{1}^{\top}\bm{P}$, which is not symmetric $\acute{\bm{C}}^{\top}\neq\acute{\bm{C}}$ nor idempotent $\acute{\bm{C}}^{2}\neq\acute{\bm{C}}$. We define the Gram matrix with the probability of data points as
$\bm{G}\coloneqq\bm{P}^{\frac{1}{2}}\acute{\bm{C}}^{\top}\bm{X}^{\top}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}=\left(\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\right)^{\top}\left(\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\right).$ (75)
[…]
With the same procedure as in Appendix B.3.2, we can show that the eigendecomposition of the Gram matrix is equivalent to PCA. The eigenvalue equation of Eq. (75) becomes
$\underbrace{\bm{P}^{\frac{1}{2}}\acute{\bm{C}}^{\top}\bm{X}^{\top}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}}_{\bm{G}}\bm{w}_{i}=\omega_{i}\bm{w}_{i}.$
Multiplying $\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}$ from the left of both sides, we obtain
$\underbrace{\bm{X}\acute{\bm{C}}\bm{P}\acute{\bm{C}}^{\top}\bm{X}^{\top}}_{\bm{\varSigma}}\left(\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\bm{w}_{i}\right)=\omega_{i}\left(\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\bm{w}_{i}\right).$
Then, the eigenvalues and the eigenvectors of the covariance matrix are
$\lambda_{i}=\omega_{i},$
$\bm{u}_{i}=\frac{1}{\sqrt{\omega_{i}}}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\bm{w}_{i},$
[…]
$\bm{u}_{i}^{\top}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}=$ $\frac{1}{\sqrt{\omega_{i}}}\bm{w}_{i}^{\top}\bm{P}^{\frac{1}{2}}\acute{\bm{C}}^{\top}\bm{X}^{\top}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}$
$=$ $\sqrt{\omega_{i}}\bm{w}_{i}^{\top}.$
Nevertheless, Eq. (75) is unavailable from double centering [Eq. (61)] of the distance matrix [Eq. (60)] by $\acute{\bm{C}}\bm{P}^{\frac{1}{2}}$ and $\bm{P}^{\frac{1}{2}}\acute{\bm{C}}^{\top}$ because the centering matrix does not erase the vector $\bm{1}$, i.e., $\bm{1}^{\top}\acute{\bm{C}}\neq\bm{0}^{\top}$ and $\acute{\bm{C}}^{\top}\bm{1}\neq\bm{0}$. Therefore, the connection between the weighted Gram matrix [Eq. (75)] and distance matrix is vague. In the following Appendix B.4, we instead introduce an alternative derivation of PCA starting from pairwise inner-product similarity, not distance.
B.4 Minimum inner-product error formulation
Here, we introduce the minimum inner product error formulation of PCA. By minimizing the inner product error, we obtain the same eigenvalues and eigenvectors as PCA. First, we introduce the function to minimize the mean squared inner product error $L$,
$L\coloneqq\left<\varepsilon^{2}\right>=\sum_{\bm{x},\bm{x}^{\prime}}p\left(\bm{x}\right)p\left(\bm{x}^{\prime}\right)\varepsilon^{2}\left(\bm{x},\bm{x}^{\prime}\right),$ (76)
where
$\varepsilon^{2}\left(\bm{x},\bm{x}^{\prime}\right)\coloneqq\frac{1}{2}\left[\bm{x}^{\top}\bm{x}^{\prime}-r\left(\bm{x}\right)r\left(\bm{x}^{\prime}\right)\right]^{2}$ (77)
is the inner product error of two data points $\bm{x}$ and $\bm{x}^{\prime}$ between the original high-dimensional and projected one-dimensional space. The direction of projection is $\bm{u}$, the unit vector, and the projected coordinate is $r\left(\bm{x}\right)\coloneqq\bm{u}^{\top}\bm{x}$. The factor $\frac{1}{2}$ is introduced for convenience of calculation and does not change the result. We would like to find the vector $\bm{u}$ that minimizes $L$ under the normalization constraint $\bm{u}^{\top}\bm{u}=1$. We introduce the Lagrangian function
$\mathcal{L}\coloneqq L+\alpha\left(1-\bm{u}^{\top}\bm{u}\right),$ (78)
where $\alpha$ is the Lagrange multiplier. The differentiation of $\varepsilon^{2}\left(\bm{x},\bm{x}^{\prime}\right)$ with respect to $\bm{u}$ is
$\frac{\partial}{\partial\bm{u}}\varepsilon^{2}\left(\bm{x},\bm{x}^{\prime}\right)$
$=$ $\left(\bm{x}^{\top}\bm{x}^{\prime}-\bm{u}^{\top}\bm{x}\bm{x}^{\prime\top}\bm{u}\right)\left(-2\bm{x}\bm{x}^{\prime\top}\bm{u}\right)$
$=$ $2\left[\left(\bm{u}^{\top}\bm{x}\right)\left(\bm{x}^{\prime\top}\bm{u}\right)\bm{x}\left(\bm{x}^{\prime\top}\bm{u}\right)-\left(\bm{x}^{\top}\bm{x}^{\prime}\right)\bm{x}\left(\bm{x}^{\prime\top}\bm{u}\right)\right]$
$=$ $2\left[\bm{x}\left(\bm{x}^{\top}\bm{u}\right)\left(\bm{u}^{\top}\bm{x}^{\prime}\right)\left(\bm{x}^{\prime\top}\bm{u}\right)-\bm{x}\left(\bm{x}^{\top}\bm{x}^{\prime}\right)\left(\bm{x}^{\prime\top}\bm{u}\right)\right]$
Thus, the differentiation of $L$ with respect to $\bm{u}$ is
$\frac{\partial}{\partial\bm{u}}L=$ $\sum_{\bm{x},\bm{x}^{\prime}}p\left(\bm{x}\right)p\left(\bm{x}^{\prime}\right)\frac{\partial}{\partial\bm{u}}\varepsilon^{2}\left(\bm{x},\bm{x}^{\prime}\right)$
$=$ $2\sum_{\bm{x},\bm{x}^{\prime}}p\left(\bm{x}\right)p\left(\bm{x}^{\prime}\right)\left(\bm{x}\bm{x}^{\top}\bm{u}\bm{u}^{\top}\bm{x}^{\prime}\bm{x}^{\prime\top}\bm{u}-\bm{x}\bm{x}^{\top}\bm{x}^{\prime}\bm{x}^{\prime\top}\bm{u}\right)$
$=$ $2\left(\acute{\bm{\varSigma}}\bm{u}\bm{u}^{\top}\acute{\bm{\varSigma}}\bm{u}-\acute{\bm{\varSigma}}\acute{\bm{\varSigma}}\bm{u}\right).$
Here, $\acute{\bm{\varSigma}}\coloneqq\sum_{\bm{x}}p\left(\bm{x}\right)\bm{x}\bm{x}^{\top}$ is the uncentered covariance matrix. Then, the method of Lagrange multipliers gives $\bm{0}=\frac{\partial}{\partial\bm{u}}\mathcal{L}=2\left(\acute{\bm{\varSigma}}\bm{u}\bm{u}^{\top}\acute{\bm{\varSigma}}\bm{u}-\acute{\bm{\varSigma}}\acute{\bm{\varSigma}}\bm{u}\right)-2\alpha\bm{u}$ or
$\acute{\bm{\varSigma}}\bm{u}\bm{u}^{\top}\acute{\bm{\varSigma}}\bm{u}-\acute{\bm{\varSigma}}\acute{\bm{\varSigma}}\bm{u}=\alpha\bm{u}.$ (79)
We will prove by contradiction that every solution $\bm{u}$ to Eq. (79) is an eigenvector of the uncentered covariance matrix $\acute{\bm{\varSigma}}$. Now, assume that $\bm{u}$ is not an eigenvector of $\acute{\bm{\varSigma}}$. Let $\acute{\bm{u}}_{i}$ be the eigenvector of $\acute{\bm{\varSigma}}$ corresponding to eigenvalue $\lambda_{i}$, i.e., $\acute{\bm{\varSigma}}\acute{\bm{u}}_{i}=\lambda_{i}\acute{\bm{u}}_{i}$. Without loss of generality, assuming that $\left\{\lambda_{i}\right\}_{i=1}^{n}$ are distinct eigenvalues for $n\geq 2$, we write the solution $\bm{u}$ as a linear combination of the corresponding eigenvectors
$\bm{u}=\sum_{i=1}^{n}\kappa_{i}\acute{\bm{u}}_{i},$ (80)
where $\kappa_{i}$ is a nonzero coefficient $\kappa_{i}\neq 0$ and $\sum_{i=1}^{n}\kappa_{i}^{2}=1$. With this expansion, $\acute{\bm{\varSigma}}\bm{u}=\sum_{i=1}^{n}\kappa_{i}\lambda_{i}\acute{\bm{u}}_{i}$, $\bm{u}^{\top}\acute{\bm{\varSigma}}\bm{u}=\sum_{i=1}^{n}\kappa_{i}^{2}\lambda_{i}$, and $\acute{\bm{\varSigma}}\acute{\bm{\varSigma}}\bm{u}=\sum_{i=1}^{n}\kappa_{i}\lambda_{i}^{2}\acute{\bm{u}}_{i}$. Plugging these into Eq. (79), we obtain
$\left(\sum_{i=1}^{n}\kappa_{i}\lambda_{i}\acute{\bm{u}}_{i}\right)\left(\sum_{i=1}^{n}\kappa_{i}^{2}\lambda_{i}\right)-\sum_{i=1}^{n}\kappa_{i}\lambda_{i}^{2}\acute{\bm{u}}_{i}=\alpha\sum_{i=1}^{n}\kappa_{i}\acute{\bm{u}}_{i},$ (81)
or
$\sum_{i=1}^{n}\left[\left(\sum_{j=1}^{n}\kappa_{j}^{2}\lambda_{j}\right)\kappa_{i}\lambda_{i}-\kappa_{i}\lambda_{i}^{2}-\alpha\kappa_{i}\right]\acute{\bm{u}}_{i}=\bm{0}.$ (82)
Since $\left\{\acute{\bm{u}}_{i}\right\}$ are linearly independent, we get a series of $n$ equations
$\alpha=\left(\sum_{j=1}^{n}\kappa_{j}^{2}\lambda_{j}-\lambda_{i}\right)\lambda_{i},$ (83)
[…]
$\lambda_{1}>\sum_{j=1}^{n}\kappa_{j}^{2}\lambda_{j}>\lambda_{n},$ (84)
because $\kappa_{i}\neq 0$ and $\sum_{i=1}^{n}\kappa_{i}^{2}=1$. Thus, the first equation of Eq. (83) becomes
$\alpha=\left(\sum_{j=1}^{n}\kappa_{j}^{2}\lambda_{j}-\lambda_{1}\right)\lambda_{1}<0$ (85)
and that of the $n$th becomes
$\alpha=\left(\sum_{j=1}^{n}\kappa_{j}^{2}\lambda_{j}-\lambda_{n}\right)\lambda_{n}\geq 0.$ (86)
These two inequalities contradict each other (i.e., $\alpha<0\leq\alpha$). Therefore, by contradiction, the solution $\bm{u}$ to Eq. (79) must be an eigenvector of the uncentered covariance matrix $\acute{\bm{\varSigma}}$, i.e., $\bm{u}=\acute{\bm{u}}_{i}$. Conversely, when $\bm{u}$ is a normalized eigenvector of $\acute{\bm{\varSigma}}$, or $\acute{\bm{\varSigma}}\bm{u}=\lambda\bm{u}$ and $\bm{u}^{\top}\bm{u}=1$, we have $\acute{\bm{\varSigma}}\bm{u}\bm{u}^{\top}\acute{\bm{\varSigma}}\bm{u}-\acute{\bm{\varSigma}}\acute{\bm{\varSigma}}\bm{u}=\acute{\bm{\varSigma}}\left(\bm{u}\bm{u}^{\top}-\bm{I}\right)\acute{\bm{\varSigma}}\bm{u}=\lambda\acute{\bm{\varSigma}}\left(\bm{u}-\bm{u}\right)=\bm{0},$ so it is a solution to Eq. (79) with $\alpha=0$.
Finally, we determine the optimal solution of the eigenvector. Let us expand the data vector, $\bm{x}=\sum_{i=1}^{N}x_{i}\acute{\bm{u}}_{i},$ with $\bm{u}=\acute{\bm{u}}_{1}.$ Then, the mean squared inner product error is
$L=$ $\frac{1}{2}\sum_{\bm{x},\bm{x}^{\prime}}p\left(\bm{x}\right)p\left(\bm{x}^{\prime}\right)\left(\sum_{i=1}^{N}x_{i}x_{i}^{\prime}-x_{1}x_{1}^{\prime}\right)^{2}$
$=$ $\frac{1}{2}\sum_{\bm{x},\bm{x}^{\prime}}p\left(\bm{x}\right)p\left(\bm{x}^{\prime}\right)\left(\sum_{i=2}^{N}x_{i}x_{i}^{\prime}\right)^{2}$
$=$ $\frac{1}{2}\sum_{\bm{x},\bm{x}^{\prime}}p\left(\bm{x}\right)p\left(\bm{x}^{\prime}\right)\sum_{i=2}^{N}\sum_{j=2}^{N}x_{i}x_{j}x_{i}^{\prime}x_{j}^{\prime}$
$=$ $\frac{1}{2}\sum_{i=2}^{N}\sum_{j=2}^{N}\left(\sum_{\bm{x}}p\left(\bm{x}\right)x_{i}x_{j}\right)\left(\sum_{\bm{x}^{\prime}}p\left(\bm{x}^{\prime}\right)x_{i}^{\prime}x_{j}^{\prime}\right)$
$=$ $\frac{1}{2}\sum_{i=2}^{N}\sum_{j=2}^{N}\left(\left<x_{i}x_{j}\right>\delta_{i,j}\right)^{2}$
$=$ $\frac{1}{2}\sum_{i=2}^{N}\lambda_{i}^{2}$
$=$ $\frac{1}{2}\left[\operatorname{tr}\left(\acute{\bm{\varSigma}}\right)-\lambda_{1}^{2}\right],$ (87)
which is minimized when $\lambda_{1}$ is the largest eigenvalue of $\acute{\bm{\varSigma}}$. When the data points and the projected coordinates are centered, $\bm{x}\leftarrow\bm{x}-\left<\bm{x}\right>$, $r\left(\bm{x}\right)\leftarrow r\left(\bm{x}\right)-\left<r\left(\bm{x}\right)\right>$ the uncentered covariance matrix becomes the covariance matrix, and the optimal solution corresponds to the first PC loading.
One can arrive at the same conclusion starting from Eq. (76) but through different calculations and proofs [123, 121, 142].
Appendix C PCA with perturbed distribution
C.1 General formulation
To expand our discussion in Sec. V, we consider the perturbation to the ideal bipolar distribution as
$p\left(\bm{s}\right)=\frac{1-\epsilon}{2}\left(\delta_{+\bm{\xi}_{1},\bm{s}}+\delta_{-\bm{\xi}_{1},\bm{s}}\right)+\epsilon\rho\left(\bm{s}\right),$ (88)
where $\bm{\xi}_{1}\in\left\{+1,-1\right\}^{N}$ is the most weighted vertex, $0\leq\epsilon\ll 1$ is the perturbation parameter, and $\rho\left(\bm{s}\right)\in\left[0,1\right]$ is any perturbation or noise distribution, satisfying $\sum_{\bm{s}}\rho\left(\bm{s}\right)=1$. We define the perturbation-mean of the vertex as $\bm{s}\right>_{\rho}\coloneqq\sum_{\bm{s}}\rho\left(\bm{s}\right)\bm{s}$. Then, the mean of the vertex becomes
[…]
and the perturbed covariance matrix becomes
[…]
$=$ $\grave{\bm{\varSigma}}+\epsilon\bm{B},$ (90)
where
$\grave{\bm{\varSigma}}\coloneqq\bm{\xi}_{1}\bm{\xi}_{1}^{\top}$ (91)
is the unperturbed covariance matrix and
[…]
is the perturbation matrix. The rank of the unperturbed covariance matrix $\grave{\bm{\varSigma}}$ is one. As we show in Sec. V, the non-zero eigenvalue is $\grave{\lambda}_{1}=\bm{\xi}_{1}^{\top}\bm{\xi}_{1}=N,$ and the corresponding eigenvector is $\grave{\bm{u}}_{1}=\frac{1}{\sqrt{N}}\bm{\xi}_{1}$. The remaining eigenvalues $\grave{\lambda}_{i\neq 1}=0$ are all zero and $(N-1)$-fold degenerate, and the eigenvectors $\left\{\grave{\bm{u}}_{i}\right\}_{i=2}^{N}$ form a complete orthonormal basis.
Now, we consider the eigenvalue equation of the perturbed covariance matrix,
$\bm{\varSigma}\bm{u}_{i}=\lambda_{i}\bm{u}_{i},$ (93)
[…]
$\bm{\varPi}\coloneqq\sum_{i=2}^{N}\grave{\bm{u}}_{i}\grave{\bm{u}}_{i}^{\top}=\bm{I}-\grave{\bm{u}}_{1}\grave{\bm{u}}_{1}^{\top}$ (94)
and solve the eigenvalue equation for the projected perturbation matrix
$\bm{\varPi}\bm{B}\bm{\varPi}^{\top}\grave{\bm{u}}_{i}=\chi_{i}\grave{\bm{u}}_{i},$ (95)
[…]
$\lambda_{1}\approx\grave{\lambda}_{1}+\epsilon\grave{\bm{u}}_{1}^{\top}\bm{B}\grave{\bm{u}}_{1}$ (96)
and
$\bm{u}_{1}\approx\grave{\bm{u}}_{1}-\epsilon\sum_{i=2}^{N}\grave{\bm{u}}_{i}\frac{\grave{\bm{u}}_{i}^{\top}\bm{B}\grave{\bm{u}}_{1}}{\grave{\lambda}_{i}-\grave{\lambda}_{1}}=\grave{\bm{u}}_{1}+\epsilon\frac{\bm{\varPi}\bm{B}\grave{\bm{u}}_{1}}{\grave{\lambda}_{1}}$ (97)
[…]
$\lambda_{i}\approx\epsilon\chi_{i}$ (98)
and
$\bm{u}_{i}\approx\grave{\bm{u}}_{i}-\epsilon\grave{\bm{u}}_{1}\frac{\grave{\bm{u}}_{1}^{\top}\bm{B}\grave{\bm{u}}_{i}}{\grave{\lambda}_{1}-\grave{\lambda}_{i}}=\grave{\bm{u}}_{i}-\epsilon\grave{\bm{u}}_{1}\frac{\grave{\bm{u}}_{1}^{\top}\bm{B}\grave{\bm{u}}_{i}}{\grave{\lambda}_{1}}.$ (99)
C.2 Quadripolar distribution
We then examine the quadripolar distribution by considering the bipolar distribution as the perturbation:
$\rho\left(\bm{s}\right)=\frac{1}{2}\left(\delta_{+\bm{\xi}_{2},\bm{s}}+\delta_{-\bm{\xi}_{2},\bm{s}}\right).$ (100)
The perturbation matrix becomes
$\bm{B}=-\bm{\xi}_{1}\bm{\xi}_{1}^{\top}+\bm{\xi}_{2}\bm{\xi}_{2}^{\top}$ (101)
because $\bm{s}\right>_{\rho}=\bm{0}$, and the projected perturbation matrix becomes
$\bm{\varPi}\bm{B}\bm{\varPi}^{\top}=$ $\left(\bm{I}-\bm{u}_{1}\bm{u}_{1}^{\top}\right)\left(-\bm{\xi}_{1}\bm{\xi}_{1}^{\top}+\bm{\xi}_{2}\bm{\xi}_{2}^{\top}\right)\left(\bm{I}-\bm{u}_{1}\bm{u}_{1}^{\top}\right)^{\top}$
$=$ $\bm{\xi}_{2}\bm{\xi}_{2}^{\top}-\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\left(\bm{\xi}_{2}\bm{\xi}_{1}^{\top}+\bm{\xi}_{1}\bm{\xi}_{2}^{\top}\right)+\left(\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\right)^{2}\bm{\xi}_{1}\bm{\xi}_{1}^{\top}$
$=$ $\left(\bm{\xi}_{2}-\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\bm{\xi}_{1}\right)\left(\bm{\xi}_{2}-\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\bm{\xi}_{1}\right)^{\top},$ (102)
which is the rank-one matrix with the nonzero eigenvalue,
$\chi_{2}=$ $\left(\bm{\xi}_{2}-\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\bm{\xi}_{1}\right)^{\top}\left(\bm{\xi}_{2}-\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\bm{\xi}_{1}\right)$
$=$ $\bm{\xi}_{2}^{\top}\bm{\xi}_{2}-2\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\bm{\xi}_{1}^{\top}\bm{\xi}_{2}+\left(\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\right)^{2}\bm{\xi}_{1}^{\top}\bm{\xi}_{1}$
$=$ $N\left[1-\left(\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\right)^{2}\right],$ (103)
and the corresponding eigenvector
$\grave{\bm{u}}_{2}=\frac{1}{\sqrt{N\left[1-\left(\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\right)^{2}\right]}}\left(\bm{\xi}_{2}-\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\bm{\xi}_{1}\right).$ (104)
With straightforward calculations, we derive the first-order approximation of the eigenvalues and eigenvectors of the first two PCs. The eigenvalue and eigenvector of PC1 are
$\lambda_{1}\approx N+\epsilon N\left[-1+\left(\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\right)^{2}\right]$ (105)
and
$\bm{u}_{1}\approx\grave{\bm{u}}_{1}-\epsilon\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\left(\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\frac{\bm{\xi}_{1}}{\sqrt{N}}-\frac{\bm{\xi}_{2}}{\sqrt{N}}\right).$ (106)
Those of PC2 are
$\lambda_{2}\approx\epsilon N\left[1-\left(\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\right)^{2}\right]$ (107)
and
$\bm{u}_{2}\approx\grave{\bm{u}}_{2}-\epsilon\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\sqrt{N\left[1-\left(\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\right)^{2}\right]}\bm{\xi}_{1}.$ (108)
C.3 Relation with Hamming distance
We define the angle $\vartheta$ between the two weighted vertices as $\vartheta\coloneqq\arccos\left(\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\right)$ and consider the projected coordinates of vertex $\bm{s}$ on PC1 and PC2 with zeroth-order approximation, obtaining
$r_{1}=\bm{u}_{1}^{\top}\bm{s}=\frac{1}{\sqrt{N}}\bm{\xi}_{1}^{\top}\bm{s}$ (109)
and
$r_{2}=\bm{u}_{2}^{\top}\bm{s}=\frac{1}{\sqrt{N}\sin\left(\vartheta\right)}\left[\bm{\xi}_{2}^{\top}\bm{s}-\bm{\xi}_{1}^{\top}\bm{s}\cos\left(\vartheta\right)\right].$ (110)
We then reveal the relationship between the projected coordinates and Hamming distance. Using Eqs. (109) and (110), we have
$\bm{\xi}_{2}^{\top}\bm{s}=$ $\sqrt{N}r_{2}\sin\left(\vartheta\right)+\bm{\xi}_{1}^{\top}\bm{s}\cos\left(\vartheta\right)$
$=$ $\sqrt{N}\left[r_{1}\cos\left(\vartheta\right)+r_{2}\sin\left(\vartheta\right)\right].$ (111)
As seen in Sec. V, the Hamming distance is a function of the inner product, $D_{\mathrm{H}}\left(\bm{\xi},\bm{s}\right)=\frac{N-\bm{\xi}^{\top}\bm{s}}{2}.$ Therefore, we express the Hamming distance as a function of the projected coordinates
$D_{\mathrm{H}}\left(\bm{\xi}_{1},\bm{s}\right)=\frac{N-\sqrt{N}r_{1}}{2}$ (112)
and
$D_{\mathrm{H}}\left(\bm{\xi}_{2},\bm{s}\right)=\frac{N-\sqrt{N}\left[r_{1}\cos\left(\vartheta\right)+r_{2}\sin\left(\vartheta\right)\right]}{2}.$ (113)
If the two weighted vertices are orthogonal, $\bm{\xi}_{1}^{\top}\bm{\xi}_{2}=0$ or $\vartheta=\frac{\uppi}{2}$, the Hamming distance from the second weighted vertex is $D_{\mathrm{H}}\left(\bm{\xi}_{2},\bm{s}\right)=\frac{N-\sqrt{N}r_{2}}{2}$ but otherwise, the Hamming distance is a function of both $r_{1}$ and $r_{2}$. The angle $\vartheta$ might not be known a priori, but it can be estimated by measuring the angle between the weighted vertices in the projected space. Indeed, the projected coordinates of the second weighted vertex on PC1 is
$\bm{u}_{1}^{\top}\bm{\xi}_{2}=\frac{1}{\sqrt{N}}\bm{\xi}_{1}^{\top}\bm{\xi}_{2}=\sqrt{N}\cos\left(\vartheta\right)$ (114)
and on PC2 is
$\bm{u}_{2}^{\top}\bm{\xi}_{2}=\frac{1}{\sqrt{N}\sin\left(\vartheta\right)}\left[\bm{\xi}_{2}^{\top}\bm{\xi}_{2}-\bm{\xi}_{1}^{\top}\bm{\xi}_{2}\cos\left(\vartheta\right)\right]=\sqrt{N}\sin\left(\vartheta\right).$ (115)
In this Appendix C, we consider the bipolar distribution [Eq. (24)] as the distribution of perturbation. In the framework of perturbation theory, one can consider any distribution as the perturbation. Thus, the extension of this discussion to any distribution is straightforward.
Appendix D Possible region of projected coordinates of hypercubic vertices
D.1 General bound of projected coordinates
In general, the orthogonally projected coordinates of hypercubic vertices are bounded by a circle with radius $\sqrt{N}$. This is because we have $\left|\bm{r}\left(\bm{s}\right)\right|\leq\sqrt{N}$ for any vertex. To show this, we consider the upper bound of $\left|\bm{r}\right|^{2}$,
$\left|\bm{r}\left(\bm{s}\right)\right|^{2}=$ $\bm{r}^{\top}\left(\bm{s}\right)\bm{r}\left(\bm{s}\right)$
$=$ $\sum_{i=1}^{2}\left(\bm{u}_{i}^{\top}\bm{s}\right)^{2}$
$\leq$ $\sum_{i=1}^{N}\left(\bm{u}_{i}^{\top}\bm{s}\right)^{2}$
$=$ $\sum_{i=1}^{N}\bm{s}^{\top}\bm{u}_{i}\bm{u}_{i}^{\top}\bm{s}$
$=$ $\bm{s}^{\top}\left(\sum_{i=1}^{N}\bm{u}_{i}\bm{u}_{i}^{\top}\right)\bm{s}$
$=$ $\bm{s}^{\top}\bm{s}$
$=$ $N$ (116)
D.2 The bound of projected coordinates with quadripolar distribution
In Sec. VI.3, we consider the projected coordinates assuming the quadripolar distribution of Eq. (100). We show the possible region of the projected coordinates in Fig. 8(b) and we derive the boundary of this region in this Appendix.
We have the triangular inequalities
$D_{\mathrm{H}}\left(\bm{\xi}_{1},\bm{\xi}_{2}\right)\leq D_{\mathrm{H}}\left(+\bm{s},\bm{\xi}_{1}\right)+D_{\mathrm{H}}\left(+\bm{s},\bm{\xi}_{2}\right)$ (117)
$D_{\mathrm{H}}\left(\bm{\xi}_{1},\bm{\xi}_{2}\right)\leq D_{\mathrm{H}}\left(-\bm{s},\bm{\xi}_{1}\right)+D_{\mathrm{H}}\left(-\bm{s},\bm{\xi}_{2}\right)$ (118)
which are similar to Eqs. (30) and (32) but different inequalities. Because weighted states are orthogonal to each other, $D_{\mathrm{H}}\left(\bm{\xi}_{1},\bm{\xi}_{2}\right)=\frac{N}{2}$ with the assumption that $N$ is even. With sufficiently large $N$, the results are qualitatively the same when $N$ is odd. Following the same procedure to derive Eq. (33), we obtain
$\frac{N}{2}\leq N-\frac{\sqrt{N}}{2}\left|r_{1}\left(\bm{s}\right)+r_{2}\left(\bm{s}\right)\right|,$ (119)
which is equivalent to Eq. (45). Similarly, we have
$D_{\mathrm{H}}\left(\bm{\xi}_{1},-\bm{\xi}_{2}\right)\leq D_{\mathrm{H}}\left(+\bm{s},\bm{\xi}_{1}\right)+D_{\mathrm{H}}\left(+\bm{s},-\bm{\xi}_{2}\right)$ (120)
$D_{\mathrm{H}}\left(\bm{\xi}_{1},-\bm{\xi}_{2}\right)\leq D_{\mathrm{H}}\left(-\bm{s},\bm{\xi}_{1}\right)+D_{\mathrm{H}}\left(-\bm{s},-\bm{\xi}_{2}\right)$ (121)
and $D_{\mathrm{H}}\left(\bm{\xi}_{1},-\bm{\xi}_{2}\right)=N-D_{\mathrm{H}}\left(\bm{\xi}_{1},\bm{\xi}_{2}\right)=\frac{N}{2}$. Repeating the same procedure to derive Eq. (33), we have
$\frac{N}{2}\leq N-\frac{\sqrt{N}}{2}\left|r_{1}\left(\bm{s}\right)-r_{2}\left(\bm{s}\right)\right|,$ (122)
which is Eq. (46). Therefore, the possible region of the projected coordinates is the square bounded by Eqs. (45) and (46), which excludes the corners of Fig. 8(b). We show the boundary of this region in Fig. 8(b).